### v1

# Toponium: the smallest bound state and simplest hadron in quantum mechanics

Abstract
We explore toponium, the smallest known quantum bound state of a top quark and its antiparticle, bound by the strong force. With a Bohr radius of $8\times 10^{-18}$ m and a lifetime of $2.5\times 10^{-25}$ s, toponium uniquely probes microphysics. Unlike all other hadrons, it is governed by ultraviolet freedom, exhibiting feeble interactions at distances much smaller than $10^{-15}$ m, rather than infrared slavery that characterizes powerful interactions at approximately $10^{-15}$ m. This distinction offers novel insights into quantum chromodynamics. Our analysis reveals a toponium signal exceeding five standard deviations in the distribution of the cross section ratio between $e^{+}e^{-}\rightarrow b\bar{b}$ and $e^{+}e^{-}\rightarrow q\bar{q}$ ($q=b,~{}c,~{}s,~{}d,~{}u$), based on 400 fb-1 (1 fb = $10^{-43}~{}{\rm m}^{2}$) of data collected at around 341 GeV, driven by quantum interference. This discovery enables a top quark mass measurement with an uncertainty reduced by a factor of ten compared to current precision levels. Detection prospects at the Circular Electron Positron Collider or the Future Circular Lepton Collider underscore their potential to revolutionize our understanding of quantum mechanics.
Introduction
The top quark ($t$), the heaviest known element particle, plays a pivotal role in Standard Model (SM) of particle physics. Its unique properties, particularly its mass, have far-reaching implications for our understanding of fundamental forces and the stability of the universe. Toponium, a bound state of a top quark and an anti-top quark ($\bar{t}$), represents the smallest quantum mechanical bound state known to date. Formed by strong force, toponium has nearly twice the mass of a single top quark [1, 2], making it about 360 times heavier than a proton and twice as massive as a gold nucleus. The Bohr radius of toponium, $8\times 10^{-18}$ m, is significantly smaller than those of hydrogen ($5.3\times 10^{-11}$ m) and proton (0.83 fm, 1 fm=$10^{-15}$ m). With an extremely short lifetime of $2.5\times 10^{-25}$ s, the study of toponium offers a unique opportunity to explore fundamental aspects of quantum mechanics, including entanglement [3, 4, 5], as well as to explore quantum gravity and extra dimensions [6, 7] at scales of $10^{-18}$ m and $10^{-25}$ s.
The top quark mass ($m_{t}$) is a crucial fundamental parameter in SM, as $m_{t}$ significantly influences our understanding of SM, particularly its vacuum stability, and provides a window into potential new physics beyond SM [8, 9, 10, 11]. Despite its importance, $m_{t}$ cannot be predicted theoretically and must be measured experimentally. The most precise measurement to date [2] yields an on-shell mass of $m_{t}^{\text{OS}}=(172.52\pm 0.14_{\rm stat.}\pm 0.30_{\rm syst.})~{}{\rm GeV}/c^{2}=(172.52\pm 0.33)~{}{\rm GeV}/c^{2}$ (GeV, gigaelectronvolts; $c$, the speed of light in vacuum, is ignored in subsequent units for simplicity). While the statistical uncertainty (${\rm stat.}$) can be reduced with longer data acquisition time, the systematic uncertainty (${\rm syst.}$) defines the ultimate precision of the measurement.
The most precise method for determining $m_{t}$ involves the near-threshold production of $t\bar{t}$ at future electron positron ($e^{+}e^{-}$) colliders, such as the Circular Electron Positron Collider (CEPC) [12, 13] and the Future Circular Lepton Collider (FCC-ee) [14]. Projections suggest that $m_{t}$ could be measured to a precision of $25-59$ MeV at CEPC [15] and $40-75$ MeV at FCC-ee [16], with statistical uncertainties as low as 9 MeV at integrated luminosities of $100-200$ fb-1 (1 fb = $10^{-43}~{}{\rm m}^{2}$) at around 343 GeV, although systematic uncertainties will dominate. These uncertainties arise from multiple sources, such as beam energy calibration, theoretical modeling, and detector performance [17, 15]. Therefore, the accurate determination of $m_{t}$ relies heavily on the development of novel method and advanced systematic control.
Despite the challenges, precise measurement of $m_{t}$ is essential for refining theoretical models such as supersymmetry and two Higgs doublet models, allowing tighter constraints on their parameter spaces [18, 19]. Furthermore, these precise $m_{t}$ determinations have far-reaching implications for cosmology and fundamental physics  [20], and the source of inflation in the universe [21], making them indispensable for current and future research. The coupling between $t$ and the Higgs boson ($H$), which generates $m_{t}$, is one of the strongest fundamental interactions, remaining significant - second only to gravity - and potentially constraining the safety of quantum gravity at energy scales far beyond the Planck scale ($10^{19}$ GeV) [22].
In addition to the on-shell top quark mass $m_{t}^{\text{OS}}$ and other scheme- and model-dependent definitions, the 1S mass of top quark, defined as $m_{t}^{\text{1S}}=m_{J_{t}}/2$ (where $J_{t}$ denotes the spin-triplet ground state of toponium) [23], constitutes the sole scheme- and model-independent measure of the top quark mass $m_{t}$ [24, 25, 26].
Furthermore, toponium distinguishes itself among hadrons because its behavior is governed by ultraviolet freedom, rather than infrared slavery of quantum chromodynamics (QCD) at distances of roughly $10^{-15}$ m [27, 28]. This unique characteristic enables toponium to be theoretically described by perturbative QCD (PQCD), thereby offering new insights into confinement within the Yang-Mills field. Given its high binding energy, the phase transition between toponium and free top quark pairs could have profound implications for the structure of the quark-gluon plasma [29], cooling rates [30], and the supercooled electroweak phase transition in the early universe [31, 32, 33].
In this study, we propose a method to discover $J_{t}$ and accurately determine $m_{t}^{\text{1S}}$ by measuring the cross section ratio:
$R_{b}=\frac{\sigma_{\rm Born}(e^{+}e^{-}\rightarrow b\bar{b})}{\sum_{q=u,d,s,c,b}\sigma_{\rm Born}(e^{+}e^{-}\rightarrow q\bar{q})}$ (1)
at $e^{+}e^{-}$ colliders near $2m_{t}^{\text{1S}}$, where $b$ is the bottom quark and $q$ is a quark for $u$$d$$s$$c$, or $b$. Future measurements of $R_{b}$ at the Z pole are expected to achieve precision levels of $4\times 10^{-5}$ at CEPC [12] and less than $6\times 10^{-5}$ at FCC-ee [34]. The relative precision of $R_{b}$ at $365$ GeV with $1.5~{}{\rm ab}^{-1}$ data at FCC-ee will be $9.6\times 10^{-4}$ [35], which results an absolute precision level of $1.4\times 10^{-4}$. The other possible methods to identify toponium can be found in Refs. [36, 37] (Further details can be found in Supplemental material.).
Modeling $J_{t}$ Interaction
We introduce the Lagrangian density $\mathcal{L}_{J_{t}}$ of the field $J_{t}$ into that of SM, $\mathcal{L}_{\text{SM}}$ [36] as follows (Supplemental material):
$\mathcal{L}_{J_{t}}$ $=$ $\sum_{f=b,e}\bar{f}\left(g^{V}_{{J_{t}}ff}\gamma_{\mu}+g^{A}_{{J_{t}}ff}\gamma_{\mu}\gamma^{5}\right)f{J_{t}}^{\mu}+\ldots,$ (2)
where the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$ are determined by non-relativistic QCD [38], $f$ represents the field of either the $b$ quark or the electron, $\gamma_{\mu}$ and $\gamma^{5}$ are the Dirac gamma matrices, and ${J_{t}}^{\mu}$ is the field of $J_{t}$.
The ${J_{t}}$ bound state is obtained by the Schrödinger equation with the Coulomb potential between $t$ and $\bar{t}$ as a lattice QCD result at an infinite quark mass [39, 40] (Supplemental material),
$V(r)=-\frac{\hbar c\lambda}{r},$ (3)
where $\hbar$ is the reduced Planck constant, the Coulombic coefficient $\lambda=0.285\pm 0.011$ [39, 40], and $r$ is the distance between $t$ and $\bar{t}$. We derive the binding energy $B_{J_{t}}$, mass $m_{J_{t}}=2m_{t}^{\text{OS}}-B_{J_{t}}=2m_{t}^{\text{1S}}$, and the square of the wave function at the origin $|\psi_{J_{t}}(0)|^{2}$:
$B_{J_{t}}$ $=$ $\frac{\lambda^{2}m_{t}^{\text{OS}}}{4}=(3.503\pm 0.270)~{}{\rm GeV},$
$m_{J_{t}}$ $=$ $(341.537\pm 0.653_{m_{t}^{\text{OS}}}\pm 0.270_{\lambda})~{}{\rm GeV}=(341.537\pm 0.707)~{}{\rm GeV},$
$|\psi_{J_{t}}(0)|^{2}$ $=$ $\frac{(\lambda m_{t}^{\text{OS}})^{3}}{8\pi}.$ (4)
The uncertainty in $m_{J_{t}}$ is mainly due to $m_{t}^{\text{OS}}$. The Bohr radius of toponium is calculated to be $8\times 10^{-18}$ m.
The $B_{J_{t}}$ and $|\psi_{J_{t}}(0)|^{2}$ have also been calculated in PQCD, but the convergence of the perturbative series up to next-to-next-to-next-to leading order (NNNLO) is still unsatisfactory (Supplemental material), requiring the consideration of higher order corrections. Including the Higgs and photon contributions, the NNNLO binding energy is 3.194 GeV, which is comparable to the $B_{J_{t}}$ in Eq.(Modeling $J_{t}$ Interaction) (Supplemental material). The estimated production cross section of spin-singlet toponium at the Large Hadron Collider (LHC) at 13 TeV, assuming identical masses and wave functions for the toponium states, is 6.24 pb (Supplemental material). This estimate agrees with the value of 6.43 pb reported in Refs. [36, 37] and the recent CMS result of $(7.1\pm 0.8)$ pb [41].
The width $\Gamma_{J_{t}}$ of $J_{t}$ is dominated by twice the width of the top quark [42] (Supplemental material):
$\Gamma_{J_{t}}=(2.603\pm 0.019)~{}{\rm GeV}.$ (5)
With the Breit-Wigner formula [43], the Feynman amplitude of the $J_{t}$ medium particle is:
${\mathcal{M}}(J_{t})\propto\frac{\lambda^{3}}{s-m_{J_{t}}^{2}+i\Gamma_{J_{t}}m_{J_{t}}},$ (6)
where $m_{J_{t}}$ and $\Gamma_{J_{t}}$ are its mass and width, $\lambda^{3}$ is part of the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$, and $\sqrt{s}$ is center of mass energy. The real part of the denominator $s-m_{J_{t}}^{2}$ goes from negative to 0 to positive as $\sqrt{s}$ goes over $m_{J_{t}}$. In the amplitudes ${\mathcal{M}}(\gamma)$ and ${\mathcal{M}}(Z)$ corresponding to the contributions of the photon (with $m_{\gamma}=\Gamma_{\gamma}=0$) and the Z boson (with $m_{Z}=91.188~{}$ GeV) in SM respectively, the sign of the term $s-m^{2}$ remains unaltered. This results in a characteristic dip-bump structure in the $R_{b}$ distribution, stemming from the fact that the production probabilities for quantum interference are proportional to the square of the sum of amplitudes, namely $|{\mathcal{M}}(J_{t})+{\mathcal{M}}(\gamma)+{\mathcal{M}}(Z)|^{2}$. This dip-bump structure in the $R_{b}$ distribution of SM with $J_{t}$ is seen in Fig. 1, while the $R_{b}$ distribution of SM without $J_{t}$ is very close to a horizontal straight line. By observing the deviation of $R_{b}$ distribution from the straight line, we can identify $J_{t}$, and a precise measurement of this deviation allows us to determine $m_{t}^{\text{1S}}$.
We choose the central values of $m_{J_{t}}$ and $\lambda$ as input, and our results will demonstrate the validity of this method within the corresponding ranges. Since $\Gamma_{J_{t}}$ can be predicted within a certain range in SM without affecting our conclusions, we fix $\Gamma_{J_{t}}$ to its central value.
Measuring $R_{b}$ and identifying the signature of $J_{t}$
The $R_{b}$, an observable which can be precisely measured in a pseudo-experiment with the CEPC scenario, is determined by [44] (Supplemental material):
$R_{b}$ $=$ $\frac{f_{\rm ISR}^{udsc}\left(N_{\text{tag}}\varepsilon_{\text{hadron}}^{udsc}-N_{\text{hadron}}\varepsilon_{\text{tag}}^{{udsc}}\right)}{N_{\text{tag}}\left(f_{\rm ISR}^{udsc}\varepsilon_{\text{hadron}}^{udsc}-f_{\rm ISR}^{b}\varepsilon^{b}_{\text{hadron}}\right)+N_{\text{hadron}}\left(f_{\rm ISR}^{b}\varepsilon_{\text{tag}}^{b}-f_{\rm ISR}^{udsc}\varepsilon_{\text{tag}}^{udsc}\right)}.$ (7)
Here, the $\varepsilon^{q}_{\text{hadron}}$ ($\varepsilon^{q}_{\text{tag}}$) is the detection efficiency (double $b$ tagged efficiency) for $q=u,~{}d,~{}s,~{}c,~{}b$ after the event selection criteria mentioned in supplementary materials, and the $\varepsilon_{\text{hadron}}^{udsc}$ ($\varepsilon_{\text{tag}}^{udsc}$) is the averaged detection efficiency (double $b$ tagged efficiency) weighted by the cross section of $e^{+}e^{-}\to q\bar{q}~{}(q=u,~{}d,~{}s,~{}c)$, where $\varepsilon_{\text{hadron}}^{{udsc}}={52.48\%}$, $\varepsilon_{\text{hadron}}^{{b}}={52.70\%}$, $\varepsilon_{\text{tag}}^{b}={46.42\%}$, and $\varepsilon_{\text{tag}}^{{udsc}}={0.04\%}$ (Supplemental material). Meanwhile, the $N_{\text{tag}}$ $(N_{\rm hadron})$ is the number of selected double $b$ tagged hadronic events (the number of hadronic events) from the experiment datasets, and can be obtained with information of the cross section, integrated luminosity, and detection efficiency at an energy point (Supplemental material). The $f^{b}_{\rm ISR}$ and $f^{udsc}_{\rm ISR}$ are the correction factors that include the initial state radiation (ISR) and energy spread (Supplementary Material). Then the $R_{b}$ value and its corresponding statistical uncertainty can be determined at an energy point.
The chi-square statistic ($\chi^{2}$) is a standard tool used to evaluate the significance of a signal or the goodness of fit of a given model. In pseudo-experiments, we perform an analysis on datasets generated by SM with $J_{t}$ model to extract $R_{b}$ value ($R_{b}^{\text{PEX}}$) at a chosen energy point ($\sqrt{s_{i}}$). The $\chi^{2}$ is then defined as follows:
$\chi^{2}_{\rm Model}=\sum_{i}{\left(\frac{R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})-R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}};m_{J_{t}},\lambda)}{\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}\right)^{2}}.$ (8)
The $\mathcal{L}_{i}$ is the integrated luminosity at $\sqrt{s_{i}}$, and ${\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}$ is the uncertainty of $R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})$. The fitting function $R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}};m_{J_{t}},\lambda)$ gives a theoretical $R_{b}$ value at an energy point, where free parameters $m_{J_{t}}$ and $\lambda$ can be varied to minimize $\chi^{2}_{\rm Model}$.
[…]
Figure 1: The blue dashed line represents $R_{b}$ distribution in SM without $J_{t}$ model, while the red solid line represents $R_{b}$ distribution in SM with $J_{t}$. Additionally, the statistical significance of three energy points is indicated. The insert shows the LO Feynman diagram for the process $e^{+}e^{-}\rightarrow q\bar{q}$, with the corresponding propagator represented by a photon, a Z boson, or a $J_{t}$.
Figure 2: Significance of the $J_{t}$ signal varies with $m_{J_{t}}$ and $\lambda$ within the $1\sigma$ range.
The main sources of systematic uncertainties in the $R_{b}$ measurements are the energy scale of $\sqrt{s}$, energy spread, detection efficiency, and backgrounds. All systematic uncertainties are more than an order of magnitude smaller than the statistical uncertainties (Supplementary Material), leading to a $J_{t}$ signal significance of $6.25\sigma$ even when considering the total systematic uncertainties.
When the true values of $m_{J_{t}}$ and $\lambda$ vary within their respective $1\sigma$ ranges, we determine the corresponding significance of the $J_{t}$ signal. This significance consistently exceeds the $5\sigma$ threshold and is shown in Fig. 2. A significance level greater than $5\sigma$ indicates that the probability of SM with $J_{t}$ is incorrect compared to SM without $J_{t}$ being correct is less than $5.7\times 10^{-7}$.
Extracting the top quark mass $m_{t}^{\text{1S}}$
The top quark mass $m_{t}^{\text{1S}}$ (along with the Coulombic coefficient $\lambda$) is extracted by minimizing $\chi^{2}_{\text{SM with}~{}J_{t}}$. The statistical uncertainty $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$ is found to be $\rm 32.2~{}MeV$, primarily determined by the second energy point $\sqrt{s_{2}}$. The systematic uncertainties, $({\Delta m_{t}^{\text{1S}}})_{\rm syst.}$, arise from several sources, including the energy scale and energy spread, detection efficiency, backgrounds, SM parameters, higher order corrections, and others (luminosity measurements and backgrounds from two-photon and $WWZ$ processes).
The uncertainty of $\sqrt{s_{2}}$ energy scale is estimated to be 2.0 MeV  at $2m_{t}^{\text{1S}}$ [45], which directly enters into the $m_{J_{t}}$ and contributes 1.0 MeV to $\Delta m_{t}^{\text{1S}}$. The center of mass energy spread is 458.4 MeV at $\sqrt{s_{2}}$ [15, 13], with an estimated uncertainty of 10%  [15, 45], contributing an additional 0.2 MeV to ${\Delta m_{t}^{\text{1S}}}$.
Experimentally, the background, as well as the detection efficiency and the weak mixing angle $\theta_{\rm W}$, can be determined primarily by the data from $Z$, $WW$, and $HZ$ peaks directly translated to the $m_{J_{t}}$ region [46]. The detection efficiency related uncertainty in $R_{b}$ is ${\varepsilon_{\text{hadron}}^{udsc}}/{\varepsilon_{\text{tag}}^{b}}$ (Supplemental material), which is consistent that in the $R_{b}$ measurement at the Z pole. After including uncertainties from Monte Carlo (MC) statistics and energy-momentum requirement, the uncertainty associated with the detection efficiency in $R_{b}$ is $19.1\times 10^{-6}$ (Supplemental material). This is about half of the total uncertainty of the $R_{b}$ measurement at the Z-pole at CEPC, which is $4\times 10^{-5}$ [12]. Consequently, this contributes 1.3 MeV to $\Delta m_{t}^{\text{1S}}$.
By comparing the background contributions estimated in theory and observed in experiment, a relative precision of $(\alpha_{s}/\pi)^{2}\sim 10^{-3}$ can be achieved, contributing 0.9 MeV to $\Delta m_{t}^{\text{1S}}$. Additionally, $\sin^{2}\theta_{\rm W}$ can attain an accuracy of $1\times 10^{-5}$ [12], thereby contributing 0.4 MeV to $\Delta m_{t}^{\text{1S}}$. The NNLO correction of $R_{b}$ yields a contribution of 0.1 MeV to $\Delta m_{t}^{\text{1S}}$. The contribution of other sources, such as luminosity and backgrounds from two-photon and $WWZ$ processes, is less than 0.2 MeV for $\Delta m_{t}^{\text{1S}}$ (Supplemental material).
Assuming these sources are independent, we combine them in quadrature, yielding a total systematic uncertainty of $({\Delta m_{t}^{\text{1S}}})_{\rm syst.}={1.9}$ MeV, as summarized in Table 1. Then we get:
$\Delta{m_{t}^{\text{1S}}}=(\pm 32.2_{\text{{stat.}}}\pm 1.9_{\text{{syst.}}})~{}{\text{MeV}}=\pm 32.3~{}{\text{MeV}}.$ (9)
The systematic uncertainty of our method is greatly improved over the predictions of previous studies [16, 15] due to simpler final states and differences in the physical quantities to be measured. The final states in Refs. [16, 15] are six-body systems from $e^{+}e^{-}\to t\bar{t}\to b\bar{b}WW$, where $W$ decays into quark pairs or lepton-neutrino pairs, while our method focuses on two-body final states $b\bar{b}$.
Table 1: Systematic uncertainties of the $m_{t}$ measurement in MeV. The “Other” category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
$\Delta m_{t}$(MeV) FCC-ee [16] CEPC [15] This work
Energy scale 3 2 1.0
Energy spread $3-5$ 0.2
Efficiency $5-45$ 1.3
Background $10-20$ $4-18$ 0.9
SM parameters ($\theta_{\rm W},~{}\alpha_{s}$) 3.2 17 $0.4$
Higher order corrections $40-45$ $9-26$ 0.1
Other 10 $<0.2$
Total syst. $40-74$ $23-58$ 1.9
The upcoming $e^{+}e^{-}$ colliders CEPC and FCC-ee aim to achieve integrated luminosities of 1000 fb-1 [13] and 1700 fb-1 [14], respectively, near the $t\bar{t}$ production threshold. With these luminosities, the systematic uncertainty on $m_{t}^{\text{1S}}$ is expected to remain at 1.9 MeV for both CEPC and FCC-ee. If the luminosities are chosen as $\mathcal{L}=(80,~{}900,~{}20)~{}{\rm fb}^{-1}$ and $\mathcal{L}=(80,~{}1600,~{}20)~{}{\rm fb}^{-1}$ for CEPC and FCC-ee at (340.076, 341.598, 342.771) GeV, respectively, our results show that, the statistical uncertainty, $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$, is estimated to be 18.6 MeV for CEPC and 13.9 MeV for FCC-ee, or 11.2 MeV when data from both colliders are combined. A detailed comparison of the top quark mass measurement methods at CEPC [15], FCC-ee [16], and this work is presented in Table 2.
Table 2: Detailed comparison of the top quark mass measurement methods: final states, signal and background cross sections, and uncertainty estimation at CEPC [15], FCC-ee [16], and this work.
Method CEPC [15] FCC-ee [16] This work
Final states $b\bar{b}f\bar{f}f\bar{f}$ $b\bar{b}f\bar{f}f\bar{f}$ $b\bar{b}$$q\bar{q}$
Measured quantity $\sigma({t\bar{t}})$ $\sigma({t\bar{t}})$ $R_{b}=\sigma({b\bar{b}})/\sigma({q\bar{q}})$
$\sigma_{\rm signal}$ $\sim 450~{}$fb [17] $\sim 450~{}$fb [17]   $\sim 900$ fb ($b\bar{b}$), $\sim$ 6000 fb ($q\bar{q}$)
[…]
$\sqrt{s}$ (GeV) $\sim 343~{}$GeV $\sim 343~{}$GeV $\sim 341~{}$GeV
Integrated luminosity 100 fb-1 200 fb-1 $400-2700$ fb-1
$(\Delta m_{t})_{\rm stat.}$ $9$ MeV $9$ MeV $11.2-32.2$ MeV
$(\Delta m_{t})_{\rm syst.}$ $23-58$ MeV   $40-74$ MeV 1.9 MeV
$(\Delta m_{t})_{\rm total}$     $25-59$ MeV     $45-75$ MeV $11.3-32.3$ MeV
Discussion and Outlook
In this study, we propose a method for both discovering toponium and precisely measuring the top quark mass $m_{t}^{\text{1S}}$ using the observable $R_{b}$ near the energy threshold $2m_{t}^{\text{1S}}$ in $e^{+}e^{-}$ collisions. The significance of the toponium signal exceeds $5\sigma$, including both statistical and systematic uncertainties with the integrated luminosity of 400 $\rm fb^{-1}$ at around 341 GeV. The $m_{t}^{\text{1S}}$ can be measured with a total uncertainty of only 32.3 MeV, including a systematic uncertainty of 1.9 MeV. This enables a top quark mass measurement with an uncertainty reduced by a factor of 10 compared to the current precision [2]. Moreover, our method improves the systematic uncertainty by more than a factor of 150 compared to the most precise current measurements [2], and by at least a factor of 12 compared to other existing methods [15, 16].
As the smallest quantum bound state and simplest hadron, toponium provides a unique opportunity to probe fundamental principles of quantum mechanics [3, 5], including particle-wave duality, entanglement, and spin dynamics, while offering novel insights into the nature of confinement in QCD. Additionally, it could play a critical role in influencing the cooling rate and supercooled electroweak phase transition in the early universe [31, 32, 33]. A precise measurement of $m_{t}$ is crucial in particle physics, as it directly impacts the stability of SM vacuum and determines whether the universe exists in a metastable state [10, 11].
However, it should be noted that while our method offers a high precision, it is sensitive to the parameter $\lambda$, which governs the coupling between $J_{t}$ and both $b\bar{b}$ and $e^{+}e^{-}$. A precise understanding of $\lambda$ is essential, and can be achieved by lattice QCD or higher order PQCD calculations.
Additionally, polarized beams or the well separation of $b$-jet and $\bar{b}$-jet in detection can reduce the integrated luminosities by an order of magnitude while maintaining the accuracy of the $m_{t}$ measurements. Alternatively, they can increase the accuracy of the $m_{t}$ measurements by a factor of three without altering the integrated luminosities. Furthermore, comparisons with the multi-body final state approach to top quark mass measurements could help to validate and complement the results obtained here.
Ultimately, the goal is to integrate this method into a broader framework for discovering new bound states and measuring the masses of fundamental particles, providing a robust and comprehensive understanding of quantum mechanics and particle physics.
Acknowledgements.
We thank Profs. Kuang-Ta Chao, Gang Li, Yan-Qing Ma, Manqi Ruan, Xiaohu Sun, and Shi-Lin Zhu for valuable and helpful discussions. This work is supported in part by National Key Research and Development Program of China under Contract No. 2020YFA0406300, and No. 2024YFA1610503, National Natural Science Foundation of China (NSFC) under contract No. 11975076, No. 12161141008, No. 12135005, No. 12075018, and No. 12405102, the Fundamental Research Funds for the Central Universities Grant No. 4007022302, and the China Postdoctoral Science Foundation under Grant No. 2024M750049.
Appendix A Supplemental material
A.1 Top quark and toponium
The most precise $m_{t}^{\text{OS}}$ measurement method is the top quark pair production at the $2m_{t}^{\text{OS}}$ threshold at the future $e^{+}e^{-}$ colliders [15, 47, 48, 49, 50, 51, 52, 17, 53, 54, 55], such as CEPC [12], Compact Linear Collider [56], FCC-ee [14], and International Linear Collider [57]. The cross sections of $e^{+}e^{-}\to t\bar{t}$ near the threshold are calculated to NNLO [53, 58, 59, 60] and NNNLO [61, 54, 62, 63, 61].
[…]
A.2 Theoretical calculation
A.2.1 Modeling $J_{t}$ interaction
The ${J_{t}}$ bound state results from the Schrödinger equation with a potential between $t$ and $\bar{t}$, which is simplified to a Coulomb plus linear potential based on lattice QCD results [39, 40]:
$V(r)=-\frac{\lambda}{r}+\sigma r,$ (10)
where $r$ is the distance between $t$ and $\bar{t}$, the Coulombic coefficient is $\lambda=0.285\pm 0.011$, and the string tension is $\sigma=0.206~{}{\rm GeV}^{2}$ [39, 40]. The Bohr radius of toponium is given by $r_{B}={2}/{(m_{t}^{\text{OS}}\lambda)}$, which evaluates to ${1}/(24.584~{}\rm{GeV})$. The $\sigma r$ term in the potential can be neglected, as its contribution to the binding energy of $J_{t}$ is $-\langle\sigma r\rangle=-{3\sigma r_{B}}/{2}=-0.013~{}{\rm GeV}$.
Additionally, the simplest bosonic string model predicts the same potential as given in Eq.(10), with $\lambda=\pi/12=0.262$ [100]. This prediction is confirmed by result from SU(3) lattice gauge theory, which suggests a slightly higher value of $\lambda\sim 0.293$ [101].
The binding energy $B_{J_{t}}$ can also be calculated in perturbative QCD (PQCD), yielding $(1.57_{\rm LO}+0.65_{\rm NLO}+0.49_{\rm NNLO}+0.25_{\rm NNNLO})~{}{\rm GeV}=2.96$ GeV at NNNLO with a renormalization scale of $\mu=32.6$ GeV for $m_{J_{t}}$ and $\mu=30$ GeV for $m_{t}^{\text{OS}}$ [102]. The difference between $m_{t}^{\text{OS}}$ and $m_{t}^{\text{1S}}$ is calculated to be $(0.428_{\rm LO}+0.368_{\rm NLO}+0.262_{\rm NNLO}+(0.174\pm 0.0004)_{\rm NNNLO})$ GeV $=(1.232\pm 0.0004)$ GeV at NNNLO with $\mu=163.508$ GeV [24]. With the center value of $m_{t}^{\text{OS}}=172.52~{}{\rm GeV}$ [2], the $\tt{QQbar\_threshold}$ package [54, 63, 103] yields:
$B_{J_{t}}^{\rm PQCD}=(1.680_{\rm LO}+0.657_{\rm NLO}+0.504_{\rm NNLO}+0.243_{\rm NNNLO})~{}{\rm GeV}=3.084~{}{\rm GeV}$ (11)
at NNNLO with $\mu=1/r_{B}=24.584$ GeV. The NkLO term is at the order of $(\alpha_{s}(\mu)/\pi)^{k}\sim 0.04^{k}$ for $k=0,1,2,3,...$. However, the convergence of the perturbative series up to NNNLO is still unsatisfactory, necessitating consideration of higher order corrections. Including contributions from the Higgs and quantum electrodynamics (QED), the binding energy is shifted by $0.110$ GeV [104], thereby bringing the NNNLO binding energy in Eq.(11) to 3.194 GeV. Such result can be compared to that of $(3.503\pm 0.270)$ GeV in Eq.(Modeling $J_{t}$ Interaction).
We introduce the Lagrangian density $\mathcal{L}_{J_{t}}$ of the field $J_{t}$ into that of SM, $\mathcal{L}_{\text{SM}}$ [36]:
$\mathcal{L}_{J_{t}}$ $=$ $\sum_{f=b,e}\bar{f}\left(g^{V}_{{J_{t}}ff}\gamma_{\mu}+g^{A}_{{J_{t}}ff}\gamma_{\mu}\gamma^{5}\right)f{J_{t}}^{\mu}+\ldots,$ (12)
The ellipse represents the kinetic energy and mass term of $J_{t}$, as well as the other interactions between $J_{t}$ and the fundamental particles. The factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$ can be determined with non-relativistic QCD (NRQCD)  [38]:
$g^{A}_{{J_{t}}ee}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\frac{2\pi\alpha m_{t}^{2}\left(3\csc^{2}\theta_{\rm W}-8\right)\sec^{2}\theta_{\rm W}}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)},$
$g^{V}_{{J_{t}}ee}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\frac{8\pi\alpha\left(m_{t}^{2}\left(6\cos\left(2\theta_{\rm W}\right)-9\right)\csc^{2}\left(2\theta_{\rm W}\right)+2m_{Z}^{2}\right)}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)},$
$g^{A}_{{J_{t}}bb}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\left(\frac{2\pi\alpha m_{t}^{2}\left(m_{t}^{2}+2m_{W}^{2}\right)\csc^{2}\theta_{\rm W}}{\sqrt{3}m_{W}^{2}\left(m_{t}^{2}+m_{W}^{2}\right)}+\frac{8\pi\alpha m_{t}^{2}\left(4\cos\left(2\theta_{\rm W}\right)-1\right)\csc^{2}\left(2\theta_{\rm W}\right)}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)}\right),$
$g^{V}_{{J_{t}}bb}=-\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\left(\frac{16\pi\alpha}{3\sqrt{3}}+\frac{2\pi\alpha m_{t}^{2}\left(m_{t}^{2}+2m_{W}^{2}\right)\csc^{2}\theta_{\rm W}}{\sqrt{3}m_{W}^{2}\left(m_{t}^{2}+m_{W}^{2}\right)}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\vspace{4cm}\left.+\frac{8\pi\alpha m_{t}^{2}\left(2\cos\left(2\theta_{\rm W}\right)+4\cos\left(4\theta_{\rm W}\right)+3\right)\csc^{2}\left(2\theta_{\rm W}\right)}{3\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)}\right).$ (13)
Here $\alpha$ is the running QED coupling constant. The Lagrangian density $\mathcal{L}_{J_{t}}$ is encoded in $\tt FeynRules$ [105]. The calculations can be done with the following packages: $\tt FeynArts$ [106], $\tt FeynCalc$ [107], $\tt MadGraph$ [108], and $\tt Whizard$ [109].
The width $\Gamma_{J_{t}}$ of $J_{t}$, dominated by twice the width of the top quark $t$, is:
$\Gamma_{J_{t}}$ $=$ $2\Gamma_{t}^{\text{OS}}\left(1-\frac{\lambda^{2}}{8}\right)+\Gamma_{\rm Annihilation}=(2.603\pm 0.019)~{}{\rm GeV},$ (14)
where $\Gamma_{t}^{\text{OS}}=(1.3148^{+0.003}_{-0.005}+0.027(m_{t}^{\text{OS}}-172.69))~{}{\rm GeV}$ [42, 110], and the coefficient $\left(1-\lambda^{2}/8\right)$ comes from corrections for the fact that bound top quarks are different from free top quarks  [111]. In addition to the corrected width of top quark, we also consider the two-body annihilation decay width of $J_{t}$, $\Gamma_{\rm Annihilation}=0.0093~{}{\rm GeV}$, based on NRQCD [38].
Since the branching fraction of $J_{t}$ to $b\bar{b}$ is relatively large, we choose the measured quantity $R_{b}$ defined in Eq.(1). The parameters are chosen as [112, 2, 113]:
$m_{t}^{\text{OS}}=(172.52\pm 0.33)~{}{\rm GeV},$ (15)
$m_{W}=(80.3692\pm 0.0133)~{}{\rm GeV},$
$m_{Z}=(91.1880\pm 0.0020)~{}{\rm GeV},$
$m_{H}=(125.20\pm 0.11)~{}{\rm GeV},$
$m_{b}(m_{J_{t}})=(2.568\pm 0.010)~{}{\rm GeV},$
$m_{c}(m_{J_{t}})=(0.575\pm 0.003)~{}{\rm GeV},$
$\alpha(m_{J_{t}})=1/(126.04\pm 0.01),$
$\alpha_{s}(m_{J_{t}})=0.09844\pm 0.00062,$
$\cos\theta_{\rm W}=m_{W}/m_{Z}.$
The $\alpha(m_{J_{t}})$, the running QCD coupling constant $\alpha_{s}(m_{J_{t}})$, the running charm quark mass $m_{c}(m_{J_{t}})$, and the running bottom quark mass $m_{b}(m_{J_{t}})$ are calculated from $\alpha(m_{Z})=1/(127.930\pm 0.008)$ and $\alpha_{s}(m_{Z})=0.1180\pm 0.0009$ [112] with $\tt{RunDec}$ package [114]. The Cabibbo-Kobayashi-Maskawa (CKM) matrix is also to be employed. However, the offset of $V_{tb}$ from 1 is much smaller than the uncertainty of $\lambda$, and the contribution of $V_{cb}$ is suppressed to less than $1/100$ of the statistical uncertainty of the $R_{b}$. Therefore, the CKM matrix can be approximated as the unit matrix. Only the $R_{b}$ uncertainty originating from $\theta_{\rm W}$ needs to be preserved, as all others are smaller than $10^{-6}$.
The production cross section of pseudoscalar toponium $\eta_{t}$ at LHC is $6.24$ pb at 13 TeV with the same mass and wave function as $J_{t}$ here and the parton distribution $\tt{NNPDF40\_lo\_as\_01180}$ [115], which is consistent with $(7.1\pm 0.8)$ pb as recently reported by CMS [41]. This result underscores the significance of future $e^{+}e^{-}$ colliders, such as CEPC and FCC-ee, in probing toponium with an unprecedented precision. By leveraging the high luminosities and clean environments of these colliders, future experiments could significantly improve upon the detection and measurement of such states.
A.2.2 Theoretical prediction of $R_{b}$ near $2m_{t}^{\rm 1S}$
The experimental cross section, denoted by the symbol $\sigma$, is given by the following expression, as presented in Refs. [43, 63]:
[…]
In this formula, $\delta_{\sqrt{s}}$ is the center of mass energy spread, which is proportional to the square of $\sqrt{s}$ [15, 13]. Due to $\delta_{\sqrt{s}}$ is $460.30.186\%\times\sqrt{2}\times 175~{}{\rm GeV}=$ 460.3 MeV for FCC-ee at 350 GeV [14], and $0.20\%\times\sqrt{2}\times 180~{}$GeV$=509.1~{}$MeV for CEPC at 360 GeV [13], we can get $\delta_{\sqrt{s}}$ is about 440 MeV for FCC-ee and 460 MeV for CEPC at around $m_{J_{t}}$. The ISR factor $F(x,\sqrt{s^{\prime}})$ is given in Refs. [116, 63, 117]. Then, we can obtain $f_{\rm ISR}$, the factor that includes the effects of energy spread and ISR.
Figure 3: The LO Feynman diagrams of the $e^{+}e^{-}\rightarrow q\bar{q}$ process comprise three diagrams: the first and second diagrams correspond to SM, while the third diagram corresponds to the $J_{t}$ contribution.
The LO Feynman diagrams of the $e^{+}e^{-}\rightarrow q\bar{q}$ are shown in Fig. 3. The LO amplitude of $e^{+}e^{-}\to q\bar{q}$ can be written as:
${\mathcal{M}}(e^{+}e^{-}\to q\bar{q})$ $=$ $\frac{g^{VV}_{q}}{s}\bar{e}.\gamma^{\mu}.e~{}\bar{q}.\gamma_{\mu}.q+\frac{g^{VA}_{q}}{s}\bar{e}.\gamma^{\mu}.e~{}\bar{q}.\gamma_{\mu}.\gamma^{5}.q$ (17)
$+~{}\frac{g^{AV}_{q}}{s}\bar{e}.\gamma^{\mu}.\gamma^{5}.e~{}\bar{q}.\gamma_{\mu}.q+\frac{g^{AA}_{q}}{s}\bar{e}.\gamma^{\mu}.\gamma^{5}.e~{}\bar{q}.\gamma_{\mu}.\gamma^{5}.q,$
where the coefficient $g^{VA}_{q}$ (similar to $g^{VV}_{q}$, $g^{AV}_{q}$, and $g^{AA}_{q}$ ) is:
$g^{VA}_{q}$ $=$ $\sum_{j=\gamma,~{}Z,~{}J_{t}}{g^{V}_{jee}}^{*}\times g^{A}_{jqq}\times\frac{s}{s-m_{j}^{2}+i\Gamma_{j}m_{j}}.$ (18)
The ${g^{V}_{jee}}^{*}$ denotes the complex conjugate of $g^{V}_{jee}$. The factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$, mass $m_{J_{t}}$, and the width $\Gamma_{J_{t}}$ are given in Subsec. A.2.1, and those for $\gamma$ and $Z$ are given in Ref.[112]. Then the cross section of $\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})$ is  [118]:
$\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})=\frac{1}{12\pi s}\left(f^{A}_{q}\left(\left|g_{q}^{AA}\right|^{2}+\left|g_{q}^{VA}\right|^{2}\right)+f^{V}_{q}\left(\left|g^{AV}_{q}\right|^{2}+\left|g_{q}^{VV}\right|^{2}\right)\right).$ (19)
The factor $f^{A}_{q}$ ($f^{V}_{q}$) includes the correction factor of the axis vector current $\bar{b}.\gamma_{\mu}.\gamma^{5}.b$ (vector current $\bar{b}.\gamma_{\mu}.b$), which is 1 at LO.
The factors $f_{A}$ and $f_{V}$ were calculated to NLO  [118], NNLO [119], and NNNLO [120, 61], and the results are described in Ref. [121]:
$f^{A}_{q}$ $=$ $f^{0}_{q}+(\delta_{qc}+\delta_{qb})f^{A}_{q,2},$
$f^{V}_{q}$ $=$ $f^{0}_{q}+(\delta_{qc}+\delta_{qb})f^{V}_{q,2},$
$f^{0}_{q}$ $=$ $1+a_{s}+\frac{3}{4}Q_{q}^{2}\frac{\alpha(\sqrt{s})}{\pi}\left(1-\frac{a_{s}}{3}\right)+a_{s}^{2}\left[\frac{365}{24}-11\zeta_{3}+n_{f}\left(-\frac{11}{12}+\frac{2}{3}\zeta_{3}\right)\right]+{\mathcal{O}}(a_{s}^{3})$
$=$ $1+a_{s}+\frac{3}{4}Q_{q}^{2}\frac{\alpha(\sqrt{s})}{\pi}\left(1-\frac{a_{s}}{3}\right)+a_{s}^{2}\left(1.9857-0.1153n_{f}\right)$
$\hskip 28.45274pt+a_{s}^{3}\left(-6.6369-1.2001n_{f}-0.005178n_{f}^{2}\right)+{\mathcal{O}}\left(a_{s}^{3}\frac{m_{b}^{2}}{s},a_{s}^{4}\right),$
$f^{A}_{q,2}$ $=$ $\frac{m_{q}^{2}}{s}\left[-6-22a_{s}+a_{s}^{2}\left(-\frac{8221}{24}+57\zeta_{2}+117\zeta_{3}+n_{f}\left(\frac{151}{12}-2\zeta_{2}-4\zeta_{3}\right)\right)+{\mathcal{O}}(a^{3}_{s})\right]$
$=$ $\frac{m_{q}^{2}}{s}\left[-6-22a_{s}+a_{s}^{2}\left(-108.140+4.48524n_{f}\right)\right]+{\mathcal{O}}\left(\frac{m_{b}^{4}}{s^{2}},a_{s}^{3}\frac{m_{b}^{2}}{s}\right),$
$f^{V}_{q,2}$ $=$ $\frac{m_{q}^{2}}{s}\left[12a_{s}+a_{s}^{2}\left(\frac{253}{2}-\frac{13}{3}n_{f}\right)+{\mathcal{O}}(a^{3}_{s})\right]$ (20)
$=$ $\frac{m_{q}^{2}}{s}\left[12a_{s}+a_{s}^{2}\left(126.500-4.33333n_{f}\right)\right]+{\mathcal{O}}\left(\frac{m_{q}^{4}}{s^{2}},a_{s}^{3}\frac{m_{q}^{2}}{s}\right),$
where the $\delta_{qb}$ ($\delta_{qc}$) indicates the need to include the bottom quark mass $m_{b}$ (the charm quark mass $m_{c}$) and $m_{u}=m_{d}=m_{s}=0$, $a_{s}=\alpha_{s}(\sqrt{s})/\pi$, $Q_{b}=Q_{s}=Q_{d}=-1/3$, and $Q_{u}=Q_{c}=Q_{t}=-1/3$. The $\zeta$ is Riemann Zeta function, and $n_{f}=5$ is the flavor number of active quarks. The terms $(\alpha_{s}/\pi)^{n}\times(m_{b}^{2}/s)^{0}$ with $n>3$ are canceled out in $R_{b}$ defined in Eq.(1), so the contribution of higher order corrections in $R_{b}$ is the ${m_{b}^{4}}/{s^{2}}\sim(\alpha_{s}/\pi)^{3}\times(m_{b}^{2}/s)<10^{-7}$. Due to $\sigma_{\rm Born}(e^{+}e^{-}\to ggg)<10^{-2}~{}$fb is very small, we dropped it from $\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})$ in Eq.(1).
In SM, the LO, NLO, NNLO, and NNNLO predictions of $R_{b}$ at $\sqrt{s}$ around $m_{J_{t}}$ are:
$R_{b,~{}{\rm LO}}^{\rm SM}$ $=$ $\left(151550.209-27.365\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
$R_{b,~{}{\rm NLO}}^{\rm SM}$ $=$ $\left(151499.646-27.362\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
$R_{b,~{}{\rm NNLO}}^{\rm SM}$ $=$ $\left(151500.422-27.363\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
$R_{b,~{}{\rm NNNLO}}^{\rm SM}$ $=$ $\left(151500.457-27.364\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6}.$ (21)
Considering that future $R_{b}$ measurement at $365$ GeV with $1.5~{}{\rm ab}^{-1}$ data at FCC-ee will have a precision of $1.4\times 10^{-4}$ [35], the NLO calculation of $R_{b}$ is sufficient, and the higher order QCD corrections of $R_{b}$ are much smaller than the systematic uncertainties and can be safely ignored. The NLO Born cross sections of $e^{+}e^{-}\to b\bar{b}$ in SM without and with $J_{t}$ are shown in Fig. 4.
Figure 4: The NLO Born cross sections of $e^{+}e^{-}\to b\bar{b}$ in SM without and with $J_{t}$.
A.2.3 Measuring $R_{b}$ near $2m_{t}^{\rm 1S}$
The $R_{b}(\sqrt{s_{i}})$ is determined by the cross sections of $\sigma(e^{+}e^{-}\to b\bar{b})$ and $\sigma(e^{+}e^{-}\to q\bar{q})$ at the selected $\sqrt{s_{i}}$:
$N_{\text{tag}}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q}\varepsilon_{\text{tag}}^{q}\sigma(e^{+}e^{-}\to q\bar{q}),$
$N_{\rm hadron}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q}\varepsilon^{q}_{\rm hadron}\sigma(e^{+}e^{-}\to q\bar{q}).$ (22)
The $N_{\text{tag}}$, $N_{\rm hadron}$, $\varepsilon^{q}_{\rm tag}$, and $\varepsilon^{q}_{\rm hadron}$ are defined after Eq.(7). Then the $R_{b}$ is defined in Eq.(7). If we ignore the small values $\varepsilon_{\text{hadron}}^{udsc}-\varepsilon_{\text{hadron}}^{b}$, $\varepsilon_{\text{tag}}^{udsc}$, and $1-{f_{\rm ISR}^{b}}/{f_{\rm ISR}^{udsc}}$ , we can get:
$R_{b}\sim\frac{N_{\text{tag}}\varepsilon_{\text{hadron}}^{udsc}}{N_{\text{hadron}}\varepsilon_{\text{tag}}^{b}}.$ (23)
This equation is used to estimate only the $R_{b}$ uncertainty associated with the detection efficiency.
A.3 Experimental simulation
In this study, MC simulations of $e^{+}e^{-}$ collisions have been performed. The physics events involving the physics process with or without $J_{t}$ are generated using $\tt Whizard$ [109], where the effects of ISR and energy spread have been taken into account. Subsequently, the generated parton-level events from $\tt Whizard$ are passed to $\tt Pythia6$ [122] for parton showering, hadronization, and the decay processes of the short-life hadrons, ensuring a realistic simulation of the final state particles. Finally, $\tt Delphes$ [123], configured with the CEPC detector conditions [12], is used to simulate the detector response, allowing for the reconstruction of observable quantities from the theoretical predictions.
A.3.1 Event generation in pseudo-experiments
The best option to identify $J_{t}$ is in the final state with double $b$-jets. However, jets originating from other quark flavors can sometimes be misidentified as $b$-jets [124] with the $b$-tagging algorithms [125], thus the contribution of other final states ($u\bar{u}$,${}d\bar{d}$$s\bar{s}$$c\bar{c}$) should be also generated. Moreover, due to limitations in detector performance, some jets with low energy or those produced at extreme angles relative to the detector’s central axis [12, 14] may go undetected, which results in an underestimation of the jet multiplicity in some events. To address this, several multi-jet backgrounds from $WW$, $ZZ$, and $ZH$ decays to four-quark final states are simulated. In order to enhance the sensitivity for $J_{t}$ signal, the event selection criteria are optimized as discussed below.
A.3.2 Event selection criteria
Jets from $q\bar{q}$ are clustered from all the well reconstructed tracks with an anti-kt algorithm [126]. To select high-quality jets, we require jet transverse momentum ($p_{t}$) greater than 15 GeV. To include the case of gluon final state radiation in $q\bar{q}$ final state, the selected events are supposed to have 2 or 3 jets. As the detector does not cover all directions, the ISR photons and jets close to the beam direction may escape detection. Thus, the effective invariant mass of the final states $\sqrt{s^{\prime}}$ needs to be considered. Consequently, $\sqrt{s^{\prime}}$ in Eq.(16) can not be obtained directly by the energies of jets [44], but can be measured well by using the information of jet angular detection with an accuracy of $0.01\%$ [127]. Based on the conservation of four-momentum of the initial and final states, $\sqrt{s^{\prime}}$ of the three-jet final states can be determined as:
$\sqrt{s^{\prime}}=\begin{cases}\sqrt{s}-\sqrt{s}\frac{\cos\theta_{1}\sin\phi_{23}+\sin\theta_{1}\left(\cot\theta_{2}\sin\phi_{31}-\cot\theta_{3}\sin\phi_{21}\right)}{\left(\cos\theta_{1}+1\right)\sin\phi_{23}+\sin\theta_{1}\left(\cot\frac{\theta_{2}}{2}\sin\phi_{31}-\cot\frac{\theta_{3}}{2}\sin\phi_{21}\right)}&\text{(Forward)}\\ \sqrt{s}-\sqrt{s}\frac{\cos\theta_{1}\sin\phi_{23}+\sin\theta_{1}\left(\cot\theta_{2}\sin\phi_{31}-\cot\theta_{3}\sin\phi_{21}\right)}{\left(\cos\theta_{1}-1\right)\sin\phi_{23}+\sin\theta_{1}\left(\tan\frac{\theta_{3}}{2}\sin\phi_{21}-\tan\frac{\theta_{2}}{2}\sin\phi_{31}\right)}&\text{(Backward)}\end{cases},$ (24)
where $\theta_{i}$ and $\phi_{i}$ are the polar and azimuthal angles of the $i-$th jet for $i=1,~{}2,~{}3$, and $\phi_{ij}=\phi_{i}-\phi_{j}$. Forward (backward) indicates that the momentum direction of the ISR photon or missing jet is the same as (opposite to) the direction of the electron beam. In the two cases of forward and backward directions, $1-{\sqrt{s^{\prime}}}/{\sqrt{s}}$ is positive in one case and negative in the other, where the positive value represents the physical solution. Similarly, in the two-jet final states, we can get:
$s^{\prime}=\frac{\sin{\theta_{1}}+\sin{\theta_{2}}-|\sin{(\theta_{1}+\theta_{2})}|}{\sin{\theta_{1}}+\sin{\theta_{2}}+|\sin{(\theta_{1}+\theta_{2})}|}\times s,$ (25)
which is consistent with the Eq.(1) of Ref. [44]. We require $\sqrt{s^{\prime}}>\sqrt{s}-10~{}{\rm GeV}$ here [128, 44, 63].
For the measurement of $R_{b}$, only events with at least two $b$-tagged jets [125] are categorized as $b\bar{b}$ events, while the remaining events are considered as light quark pair processes. For the selected signal events, we require the invariant mass of jet pairs $M_{bb}$ or $M_{j_{1}j_{2}}$ greater than 150 GeV to suppress the backgrounds from W/Z hadronic decays, where ${j_{1}j_{2}}$ denote the two jets with the highest and the second-highest energies. Events passing all of the criteria are recognized as signal events, and the detection efficiencies of different processes in the three energy points are shown in Table 3.
Table 3: The experimental cross sections $\sigma(\sqrt{s})$ in fb, double $b$ tagged efficiencies $\varepsilon_{\text{tag}}(\%)$, and detection efficiencies $\varepsilon_{\text{hadron}}(\%)$ for different processes in $R_{b}$ measurements. Here $jj$ is $u\bar{u}$,${}d\bar{d}$$s\bar{s}$$c\bar{c}$, or $gg$.
Process $\sigma(340.076)$ (fb)    $\sigma(341.598)$ (fb)    $\sigma(342.771)$ (fb)   $\varepsilon_{\text{tag}}(\%)$   $\varepsilon_{\text{hadron}}(\%)$
$b\bar{b}$ 866.99 880.50 896.87 46.42 52.70
$jj$ 4977.58 4945.68 4927.87 0.04 52.48
$ZZ\to b\bar{b}b\bar{b}$ 16.17 16.01 15.86 0.31 0.31
$ZH\to b\bar{b}b\bar{b}$ 12.14 11.88 11.70 0.31 0.31
$ZZ\to b\bar{b}jj$ 116.42 115.11 114.19 1.64 2.21
$ZH\to b\bar{b}jj$ 46.06 45.11 44.42 1.64 2.21
$ZZ\to jjjj$ 209.21 207.11 205.38 0.00 5.19
$ZH\to jjjj$ 8.60 8.42 8.29 0.00 5.19
$WW\to jjjj$ 5263.55 5202.29 5157.25 0.00 5.19
The cross section $\sigma(e^{+}e^{-}\to b\bar{b})$ is about 900 fb at around 340 GeV as shown in Fig. 4, and the $\sum_{q=u,d,s,c,b}\sigma(e^{+}e^{-}\to q\bar{q})$ is about $6000$ fb at around 340 GeV. After our event selections, around 52% of the signal events passed the event selections, while nearly ${95}\%$ of the $WW/ZZ/ZH$ background events are rejected. Particularly, more than ${99}\%$ of the four-quark backgrounds are rejected in double $b$-tagged category. Once the efficiencies of all the involved processes have been established, the $R_{b}$ values at the three energy points can be obtained together with information of integrated luminosities, the cross sections, and the estimated numbers of signal and background events.
A.4 Systematic uncertainties
The systematic uncertainties, affecting the accuracy of $R_{b}$ measurements, arise from accelerator effects (e.g., $\sqrt{s}$ energy scale and energy spread), detector efficiencies (e.g., MC statistics, $\sqrt{s^{\prime}}$ and $p_{t}$ requirement), SM parameters ($\theta_{\rm W}$, $\alpha_{s}$, etc.), theoretical considerations (e.g., higher order corrections and ISR), and others (e.g., luminosity measurements and backgrounds from two-photon and $WWZ$ processes). All the systematic uncertainties are summarized in Table 4. Taking the $R_{b}$ measurement at the second energy point as an example, systematic uncertainties are discussed below in detail.
The uncertainties related to accelerator include the $\sqrt{s}$ energy scale and energy spread. The uncertainty in the energy scale of $\sqrt{s_{2}}$ is estimated to be 2 MeV, the same as that at FCC-ee at 365 GeV [45]. The uncertainty of the energy scale contributes $14.6\times 10^{-6}$ to $\Delta R_{b}$. The energy spread is 458.4 MeV for CEPC at the second energy point. A 10% uncertainty [15, 45] in the energy spread for CEPC or FCC-ee contributes an additional $2.7\times 10^{-6}$ to $\Delta R_{b}$. The uncertainties associated with the accelerator are different at the three energy points, consistent with the $R_{b}$ distribution in Fig. 1. Notably, the first and third energy points corresponding to the minimum and maximum values of $R_{b}$ are sensitive to the energy spread, while the second energy point with the largest differential of $R_{b}$ over $m_{J_{t}}$ (very close to $\sqrt{s_{2}}$) is sensitive to the energy scale.
The uncertainty associated with the detection efficiency in $R_{b}$ is ${\varepsilon_{\text{hadron}}^{udsc}}/{\varepsilon_{\text{tag}}^{b}}$, including MC statistics and the energy-momentum requirement, and is consistent with that in the $R_{b}$ measurement at the Z pole. In addition, the sizes of simulated events can be increased to achieve a statistical error less than $10^{-4}$, thus contributing $15.4\times 10^{-6}$ to $\Delta R_{b}$. We require $p_{t}>15~{}$ GeV and $\sqrt{s^{\prime}}>\sqrt{s}-10~{}$ GeV to select signal candidates, where the resolutions of $p_{t}$ and $\sqrt{s^{\prime}}$ also affect the detection efficiency. The jet energy and angular scale of the CEPC detector can be controlled within 0.5% and 0.01% [127]. Thus, the precisions of $p_{t}$ and $\sqrt{s^{\prime}}$ can be controlled to within 0.10 GeV and 0.34 GeV, respectively, resulting in $\Delta R_{b}$ values of $9.1\times 10^{-6}$ and $6.7\times 10^{-6}$.
The dominant backgrounds originate from four-fermion final states, such as $WW$, $ZZ$, and $ZH$ related processes. These processes are well understood theoretically and experimentally, and about 95% of such events are rejected by selection criteria. After collecting large data at the $Z/WW/ZH$ peaks at CEPC and FCC-ee, and taking into account the NNLO calculation, the relative uncertainty in these backgrounds will be less than $(\alpha_{s}/\pi)^{2}\sim 0.001$, thereby contributing $13.9\times 10^{-6}$ to $\Delta R_{b}$.
The uncertainty in $\sin^{2}\theta_{\rm W}$, which can be achieved at $1\times 10^{-5}$ at CEPC [12, 45], contributes $6.0\times 10^{-6}$ to ${\Delta R_{b}}$. The uncertainty of $m_{b}$ in Eq.(15) contributes $0.2\times 10^{-6}$ to ${\Delta R_{b}}$. The uncertainty from $\alpha_{s}$ can be included in the higher order corrections. As shown in Eq.(A.2.2), the NNLO corrections will increase $R_{b}$ by $1.2\times 10^{-6}$. The uncertainty associated with ISR is canceled in $R_{b}$ measurement [43].
The other sources include luminosity measurements and backgrounds from two-photon and $WWZ$ processes. Their contributions to ${\Delta R_{b}}$ are small, with a combined estimation of $<3.0\times 10^{-6}$.
Assuming all sources of systematic uncertainties in $R_{b}$ are independent, their contributions are combined in quadrature to result in a total systematic uncertainty of $R_{b}$ at the second energy point:
$({\Delta R_{b}})_{\text{syst.}}=28.6\times 10^{-6}.$ (26)
As shown in Fig. 1, the difference in $R_{b}$ between the first and second points is $(5567.6\pm 1002.1_{\rm stat.})\times 10^{-6}$, compared to SM theoretical prediction of $-41.9\times 10^{-6}$. The total systematic uncertainty in the energy scale and energy spread is ±$\pm 39.0\times 10^{-6}$, which means that the difference in $R_{b}$ is $5.6\sigma$. Other systematic uncertainties in the detection efficiency, backgrounds, etc. do not change this difference.
Table 4: Uncertainties of $R_{b}^{\rm PEX}$ are given in a unit of $10^{-6}$. The “Other” category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes. As a comparison, the pseudo-experiment $R_{b}^{\rm PEX}$ and SM prediction $R_{b}^{\rm SM}$ (in a unit of $10^{-6}$) are also provided here.
[…]
Energy scale 0.5 14.6 0.1
Energy spread 36.0 2.7 58.9
MC statistics 14.8 15.4 15.8
$p_{t}$ requirement 8.9 9.1 9.3
$\sqrt{s^{\prime}}$ requirement 6.4 6.7 7.0
Background 13.5 13.9 14.2
SM parameters 5.0 6.0 6.8
Higher order corrections 1.2 1.2 1.2
Other $<3.0$ $<3.0$ $<3.0$
${\rm Sum~{}of~{}syst.}$ 43.0 28.6 64.0
${\rm Stat.}$ 884.4 471.2 1859.3
${\rm Total}$ 885.5 472.1 1860.4
$R_{b}^{{\rm PEX}}(\sqrt{s_{i}})$     146970.0     152537.6     156589.0
$R_{b}^{{\rm SM}}(\sqrt{s_{i}})$ 151539.9 151498.0 151466.1

### v2

@@
 With a Bohr radius of $8\times 10^{-18}$ m and a lifetime of $2.5\times 10^{-25}$ s, toponium uniquely probes microphysics.
-Unlike all other hadrons, it is governed by ultraviolet freedom, exhibiting feeble interactions at distances much smaller than $10^{-15}$ m, rather than infrared slavery that characterizes powerful interactions at approximately $10^{-15}$ m.
+Unlike all other hadrons, it is governed by ultraviolet freedom.
 This distinction offers novel insights into quantum chromodynamics.
-Our analysis reveals a toponium signal exceeding five standard deviations in the distribution of the cross section ratio between $e^{+}e^{-}\rightarrow b\bar{b}$ and $e^{+}e^{-}\rightarrow q\bar{q}$ ($q=b,~{}c,~{}s,~{}d,~{}u$), based on 400 fb-1 (1 fb = $10^{-43}~{}{\rm m}^{2}$) of data collected at around 341 GeV, driven by quantum interference.
+Our analysis reveals a toponium signal exceeding $5\sigma$ in the distribution of the cross section ratio between $e^{+}e^{-}\rightarrow b\bar{b}$ and $e^{+}e^{-}\rightarrow q\bar{q}$ ($q=b,c,s,d,u$), based on 400 fb-1 of data collected at $\sqrt{s}\approx 341~{}{\rm GeV}$.
 This discovery enables a top quark mass measurement with an uncertainty reduced by a factor of ten compared to current precision levels.
-Detection prospects at the Circular Electron Positron Collider or the Future Circular Lepton Collider underscore their potential to revolutionize our understanding of quantum mechanics.
-Introduction
-The top quark ($t$), the heaviest known element particle, plays a pivotal role in Standard Model (SM) of particle physics.
+Moreover, this method improves the systematic uncertainty by at least a factor of 12 compared to any other possible methods.
+Introduction—The top quark ($t$), the heaviest known element particle, plays a pivotal role in Standard Model (SM) of particle physics.
 Its unique properties, particularly its mass, have far-reaching implications for our understanding of fundamental forces and the stability of the universe.
 Toponium, a bound state of a top quark and an anti-top quark ($\bar{t}$), represents the smallest quantum mechanical bound state known to date.
-Formed by strong force, toponium has nearly twice the mass of a single top quark [1, 2], making it about 360 times heavier than a proton and twice as massive as a gold nucleus.
-The Bohr radius of toponium, $8\times 10^{-18}$ m, is significantly smaller than those of hydrogen ($5.3\times 10^{-11}$ m) and proton (0.83 fm, 1 fm=$10^{-15}$ m).
-With an extremely short lifetime of $2.5\times 10^{-25}$ s, the study of toponium offers a unique opportunity to explore fundamental aspects of quantum mechanics, including entanglement [3, 4, 5], as well as to explore quantum gravity and extra dimensions [6, 7] at scales of $10^{-18}$ m and $10^{-25}$ s.
+Formed by strong force, toponium has nearly twice the mass of a single top quark [1, 2].
+With the smallest Bohr radius of $8\times 10^{-18}$ m and an extremely short lifetime of $2.5\times 10^{-25}$ s, the study of toponium offers a unique opportunity to explore fundamental aspects of quantum mechanics, including entanglement [3, 4, 5, 6], as well as to explore quantum gravity and extra dimensions [7, 8] at scales of $10^{-18}$ m and $10^{-25}$ s.
 The top quark mass ($m_{t}$) is a crucial fundamental parameter in SM, as $m_{t}$ significantly influences our understanding of SM, particularly its vacuum stability, and provides a window into potential new physics beyond SM [9, 10, 11, 12].
 Despite its importance, $m_{t}$ cannot be predicted theoretically and must be measured experimentally.
-The most precise measurement to date [2] yields an on-shell mass of $m_{t}^{\text{OS}}=(172.52\pm 0.14_{\rm stat.}\pm 0.30_{\rm syst.})~{}{\rm GeV}/c^{2}=(172.52\pm 0.33)~{}{\rm GeV}/c^{2}$ (GeV, gigaelectronvolts; $c$, the speed of light in vacuum, is ignored in subsequent units for simplicity).
+The most precise measurement to date [2] yields an on-shell mass of $m_{t}^{\text{OS}}=(172.52\pm 0.14_{\rm stat.}\pm 0.30_{\rm syst.})~{}{\rm GeV}=(172.52\pm 0.33)~{}{\rm GeV}$.
 While the statistical uncertainty (${\rm stat.}$) can be reduced with longer data acquisition time, the systematic uncertainty (${\rm syst.}$) defines the ultimate precision of the measurement.
 The most precise method for determining $m_{t}$ involves the near-threshold production of $t\bar{t}$ at future electron positron ($e^{+}e^{-}$) colliders, such as the Circular Electron Positron Collider (CEPC) [13, 14] and the Future Circular Lepton Collider (FCC-ee) [15].
-Projections suggest that $m_{t}$ could be measured to a precision of $25-59$ MeV at CEPC [15] and $40-75$ MeV at FCC-ee [16], with statistical uncertainties as low as 9 MeV at integrated luminosities of $100-200$ fb-1 (1 fb = $10^{-43}~{}{\rm m}^{2}$) at around 343 GeV, although systematic uncertainties will dominate.
+Projections suggest that $m_{t}$ could be measured to a precision of $25-59$ MeV at CEPC [16] and $40-75$ MeV at FCC-ee [17], with statistical uncertainties as low as 9 MeV at integrated luminosities of $100-200$ fb-1 around 343 GeV, although systematic uncertainties will dominate.
 These uncertainties arise from multiple sources, such as beam energy calibration, theoretical modeling, and detector performance [18, 16].
@@
 In addition to the on-shell top quark mass $m_{t}^{\text{OS}}$ and other scheme- and model-dependent definitions, the 1S mass of top quark, defined as $m_{t}^{\text{1S}}=m_{J_{t}}/2$ (where $J_{t}$ denotes the spin-triplet ground state of toponium) [24], constitutes the sole scheme- and model-independent measure of the top quark mass $m_{t}$ [25, 26, 27].
-Furthermore, toponium distinguishes itself among hadrons because its behavior is governed by ultraviolet freedom, rather than infrared slavery of quantum chromodynamics (QCD) at distances of roughly $10^{-15}$ m [27, 28].
-This unique characteristic enables toponium to be theoretically described by perturbative QCD (PQCD), thereby offering new insights into confinement within the Yang-Mills field.
+Furthermore, toponium distinguishes itself among hadrons because its behavior is governed by ultraviolet freedom, rather than infrared slavery of quantum chromodynamics (QCD) at distances of roughly $200~{}{\rm MeV}$ [28, 29].
+This unique characteristic enables toponium to be theoretically described by perturbative QCD (PQCD), thereby offering new insights into confinement within QCD [28].
 Given its high binding energy, the phase transition between toponium and free top quark pairs could have profound implications for the structure of the quark-gluon plasma [30], cooling rates [31], and the supercooled electroweak phase transition in the early universe [32, 33, 34].
-In this study, we propose a method to discover $J_{t}$ and accurately determine $m_{t}^{\text{1S}}$ by measuring the cross section ratio:
+In this Letter, we propose a method to discover $J_{t}$ and accurately determine $m_{t}^{\text{1S}}$ by measuring the cross section ratio:
 $R_{b}=\frac{\sigma_{\rm Born}(e^{+}e^{-}\rightarrow b\bar{b})}{\sum_{q=u,d,s,c,b}\sigma_{\rm Born}(e^{+}e^{-}\rightarrow q\bar{q})}$ (1)
@@
 Future measurements of $R_{b}$ at the $Z$ pole are expected to achieve precision levels of $4\times 10^{-5}$ at CEPC [13] and less than $6\times 10^{-5}$ at FCC-ee [35].
-The relative precision of $R_{b}$ at $365$ GeV with $1.5~{}{\rm ab}^{-1}$ data at FCC-ee will be $9.6\times 10^{-4}$ [35], which results an absolute precision level of $1.4\times 10^{-4}$.
+The relative precision of $R_{b}$ at $365$ GeV with $1.5~{}{\rm ab}^{-1}$ data will be $9.6\times 10^{-4}$ at FCC-ee [36], which results an absolute precision level of $1.4\times 10^{-4}$.
 The other possible methods to identify toponium can be found in Refs.
-[36, 37] (Further details can be found in Supplemental material.).
-Modeling $J_{t}$ Interaction
-We introduce the Lagrangian density $\mathcal{L}_{J_{t}}$ of the field $J_{t}$ into that of SM, $\mathcal{L}_{\text{SM}}$ [36] as follows (Supplemental material):
+[37, 38, 39] and related papers.
+Modeling $J_{t}$ Interaction—We introduce the Lagrangian density $\mathcal{L}_{J_{t}}$ of the field $J_{t}$ into that of SM, $\mathcal{L}_{\text{SM}}$ [37] as follows:
 $\mathcal{L}_{J_{t}}$ $=$ $\sum_{f=b,e}\bar{f}\left(g^{V}_{{J_{t}}ff}\gamma_{\mu}+g^{A}_{{J_{t}}ff}\gamma_{\mu}\gamma^{5}\right)f{J_{t}}^{\mu}+\ldots,$ (2)
-where the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$ are determined by non-relativistic QCD [38], $f$ represents the field of either the $b$ quark or the electron, $\gamma_{\mu}$ and $\gamma^{5}$ are the Dirac gamma matrices, and ${J_{t}}^{\mu}$ is the field of $J_{t}$.
-The ${J_{t}}$ bound state is obtained by the Schrödinger equation with the Coulomb potential between $t$ and $\bar{t}$ as a lattice QCD result at an infinite quark mass [39, 40] (Supplemental material),
-$V(r)=-\frac{\hbar c\lambda}{r},$ (3)
-where $\hbar$ is the reduced Planck constant, the Coulombic coefficient $\lambda=0.285\pm 0.011$ [39, 40], and $r$ is the distance between $t$ and $\bar{t}$.
+where the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$ are determined by non-relativistic QCD (NRQCD) [40], $f$ represents the field of either the $b$ quark or the electron, and ${J_{t}}^{\mu}$ is the field of $J_{t}$.
+The ${J_{t}}$ bound state is obtained by the Schrödinger equation with the Coulomb potential between $t$ and $\bar{t}$ as a lattice QCD result at an infinite quark mass [41, 42],
+$V(r)=-\frac{\lambda}{r}.$ (3)
+The Coulombic coefficient $\lambda=0.285\pm 0.011$ [41, 42], and $r$ is the distance between $t$ and $\bar{t}$.
 We derive the binding energy $B_{J_{t}}$, mass $m_{J_{t}}=2m_{t}^{\text{OS}}-B_{J_{t}}=2m_{t}^{\text{1S}}$, and the square of the wave function at the origin $|\psi_{J_{t}}(0)|^{2}$:
 $B_{J_{t}}$ $=$ $\frac{\lambda^{2}m_{t}^{\text{OS}}}{4}=(3.503\pm 0.270)~{}{\rm GeV},$
-$m_{J_{t}}$ $=$ $(341.537\pm 0.653_{m_{t}^{\text{OS}}}\pm 0.270_{\lambda})~{}{\rm GeV}=(341.537\pm 0.707)~{}{\rm GeV},$
+$m_{J_{t}}$ $=$ $(341.537\pm 0.653_{m_{t}^{\text{OS}}}\pm 0.270_{\lambda})~{}{\rm GeV}$
+$=$ $(341.537\pm 0.707)~{}{\rm GeV},$
 $|\psi_{J_{t}}(0)|^{2}$ $=$ $\frac{(\lambda m_{t}^{\text{OS}})^{3}}{8\pi}.$ (4)
 The uncertainty in $m_{J_{t}}$ is mainly due to $m_{t}^{\text{OS}}$. The Bohr radius of toponium is calculated to be $8\times 10^{-18}$ m.
-The $B_{J_{t}}$ and $|\psi_{J_{t}}(0)|^{2}$ have also been calculated in PQCD, but the convergence of the perturbative series up to next-to-next-to-next-to leading order (NNNLO) is still unsatisfactory (Supplemental material), requiring the consideration of higher order corrections.
-Including the Higgs and photon contributions, the NNNLO binding energy is 3.194 GeV, which is comparable to the $B_{J_{t}}$ in Eq.(Modeling $J_{t}$ Interaction) (Supplemental material).
-The estimated production cross section of spin-singlet toponium at the Large Hadron Collider (LHC) at 13 TeV, assuming identical masses and wave functions for the toponium states, is 6.24 pb (Supplemental material).
+The $B_{J_{t}}$ and $|\psi_{J_{t}}(0)|^{2}$ have also been calculated in PQCD, but the convergence of the perturbative series up to next-to-next-to-next-to-leading order (NNNLO) is still unsatisfactory (End Matter), requiring the consideration of higher order corrections.
+Including the Higgs and photon contributions [43], the NNNLO binding energy is 3.194 GeV, which is comparable to the $B_{J_{t}}$ in Eq.(Toponium: the smallest bound state and simplest hadron in quantum mechanics).
+The estimated production cross section of spin-singlet toponium at the Large Hadron Collider (LHC) at 13 TeV, assuming identical masses and wave functions for the toponium states, is 6.24 pb.
 This estimate agrees with the value of 6.43 pb reported in Refs.
 [37, 38] and the recent CMS result of $(7.1\pm 0.8)$ pb [44].
-The width $\Gamma_{J_{t}}$ of $J_{t}$ is dominated by twice the width of the top quark [42] (Supplemental material):
-$\Gamma_{J_{t}}=(2.603\pm 0.019)~{}{\rm GeV}.$ (5)
+The width $\Gamma_{J_{t}}$ of $J_{t}$ is dominated by twice the width of the top quark $\Gamma_{t}$ [45, 46]:
+$\Gamma_{J_{t}}$ $=$ $2\Gamma_{t}\left(1-\frac{\lambda^{2}}{8}\right)+\Gamma_{\rm Anni}=(2.603\pm 0.019)~{}{\rm GeV},$ (5)
+where the coefficient $\left(1-\lambda^{2}/8\right)$ comes from corrections for the fact that bound top quarks are different from free top quarks [47].
+The annihilation decay width of $J_{t}$, $\Gamma_{\rm Anni}=0.0093~{}{\rm GeV}$, is calculated using NRQCD [40].
 With the Breit-Wigner formula [48], the Feynman amplitude of the $J_{t}$ medium particle is:
 ${\mathcal{M}}(J_{t})\propto\frac{\lambda^{3}}{s-m_{J_{t}}^{2}+i\Gamma_{J_{t}}m_{J_{t}}},$ (6)
-where $m_{J_{t}}$ and $\Gamma_{J_{t}}$ are its mass and width, $\lambda^{3}$ is part of the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$, and $\sqrt{s}$ is center of mass energy.
+where $m_{J_{t}}$ and $\Gamma_{J_{t}}$ are its mass and width, $\lambda^{3}$ is part of the coupling factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$, and $\sqrt{s}$ is the center of mass energy.
 The real part of the denominator $s-m_{J_{t}}^{2}$ goes from negative to 0 to positive as $\sqrt{s}$ goes over $m_{J_{t}}$.
@@
 Since $\Gamma_{J_{t}}$ can be predicted within a certain range in SM without affecting our conclusions, we fix $\Gamma_{J_{t}}$ to its central value.
-Measuring $R_{b}$ and identifying the signature of $J_{t}$
-The $R_{b}$, an observable which can be precisely measured in a pseudo-experiment with the CEPC scenario, is determined by [44] (Supplemental material):
-$R_{b}$ $=$ $\frac{f_{\rm ISR}^{udsc}\left(N_{\text{tag}}\varepsilon_{\text{hadron}}^{udsc}-N_{\text{hadron}}\varepsilon_{\text{tag}}^{{udsc}}\right)}{N_{\text{tag}}\left(f_{\rm ISR}^{udsc}\varepsilon_{\text{hadron}}^{udsc}-f_{\rm ISR}^{b}\varepsilon^{b}_{\text{hadron}}\right)+N_{\text{hadron}}\left(f_{\rm ISR}^{b}\varepsilon_{\text{tag}}^{b}-f_{\rm ISR}^{udsc}\varepsilon_{\text{tag}}^{udsc}\right)}.$ (7)
-Here, the $\varepsilon^{q}_{\text{hadron}}$ ($\varepsilon^{q}_{\text{tag}}$) is the detection efficiency (double $b$ tagged efficiency) for $q=u,~{}d,~{}s,~{}c,~{}b$ after the event selection criteria mentioned in supplementary materials, and the $\varepsilon_{\text{hadron}}^{udsc}$ ($\varepsilon_{\text{tag}}^{udsc}$) is the averaged detection efficiency (double $b$ tagged efficiency) weighted by the cross section of $e^{+}e^{-}\to q\bar{q}~{}(q=u,~{}d,~{}s,~{}c)$, where $\varepsilon_{\text{hadron}}^{{udsc}}={52.48\%}$, $\varepsilon_{\text{hadron}}^{{b}}={52.70\%}$, $\varepsilon_{\text{tag}}^{b}={46.42\%}$, and $\varepsilon_{\text{tag}}^{{udsc}}={0.04\%}$ (Supplemental material).
-Meanwhile, the $N_{\text{tag}}$ $(N_{\rm hadron})$ is the number of selected double $b$ tagged hadronic events (the number of hadronic events) from the experiment datasets, and can be obtained with information of the cross section, integrated luminosity, and detection efficiency at an energy point (Supplemental material).
-The $f^{b}_{\rm ISR}$ and $f^{udsc}_{\rm ISR}$ are the correction factors that include the initial state radiation (ISR) and energy spread (Supplementary Material).
-Then the $R_{b}$ value and its corresponding statistical uncertainty can be determined at an energy point.
+Measuring $R_{b}$—The $R_{b}$, an observable which can be precisely measured in a pseudo-experiment with the CEPC scenario, is obtained by [49] (End Matter):
+$R_{b}=\frac{N_{\text{tag}}F_{\text{hadron}}^{udsc}-N_{\text{hadron}}F_{\text{tag}}^{{udsc}}}{N_{\text{tag}}(F_{\text{hadron}}^{udsc}-F^{b}_{\text{hadron}})+N_{\text{hadron}}(F_{\text{tag}}^{b}-F_{\text{tag}}^{udsc})}.$ (7)
+Here, $F=\varepsilon f_{\rm ISR}$, the detection efficiency for $b$ quarks after event selection criteria is denoted by $\varepsilon^{b}_{\text{hadron}}$ (for hadronic events) and $\varepsilon^{b}_{\text{tag}}$ (for double $b$-tagged events).
+Similarly, the averaged detection efficiency for light quarks ($q=u,d,s,c$) is denoted by $\varepsilon_{\text{hadron}}^{udsc}$ (for hadronic events) and $\varepsilon_{\text{tag}}^{udsc}$ (for double $b$-tagged events), weighted by the cross section of $e^{+}e^{-}\to q\bar{q}$.
+The values are $\varepsilon_{\text{hadron}}^{udsc}=52.48\%$, $\varepsilon_{\text{hadron}}^{b}=52.70\%$, $\varepsilon_{\text{tag}}^{b}=46.42\%$, and $\varepsilon_{\text{tag}}^{udsc}=0.04\%$ (End Matter).
+The correction factors $f^{udsc}_{\rm ISR}$ and $f^{b}_{\rm ISR}$ are used to model the effects of initial state radiation (ISR) and energy spread.
+Meanwhile, $N_{\text{tag}}$ $(N_{\rm hadron})$ is the number of selected double $b$-tagged hadronic events (the number of hadronic events) from the experiment datasets, and can be obtained with information of the cross section, integrated luminosity, and detection efficiency at an energy point.
+If we ignore the small values $\varepsilon_{\text{hadron}}^{udsc}-\varepsilon_{\text{hadron}}^{b}$, $\varepsilon_{\text{tag}}^{udsc}$, and $1-{f_{\rm ISR}^{b}}/{f_{\rm ISR}^{udsc}}$, we can get $R_{b}\sim(N_{\text{tag}}\varepsilon_{\text{hadron}}^{udsc})/(N_{\text{hadron}}\varepsilon_{\text{tag}}^{b})$.
+Then the $R_{b}$ value and its corresponding statistical uncertainty can be determined at an energy point (End Matter).
 The chi-square statistic ($\chi^{2}$) is a standard tool used to evaluate the significance of a signal or the goodness of fit of a given model.
@@
 The $\chi^{2}$ is then defined as follows:
-$\chi^{2}_{\rm Model}=\sum_{i}{\left(\frac{R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})-R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}};m_{J_{t}},\lambda)}{\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}\right)^{2}}.$ (8)
+$\chi^{2}_{\rm Model}=\sum_{i}{\left(\frac{R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})-R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}})}{\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}\right)^{2}}.$ (8)
 The $\mathcal{L}_{i}$ is the integrated luminosity at $\sqrt{s_{i}}$, and ${\Delta R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})}$ is the uncertainty of $R_{b}^{\text{PEX}}(\mathcal{L}_{i},\sqrt{s_{i}})$.
-The fitting function $R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}};m_{J_{t}},\lambda)$ gives a theoretical $R_{b}$ value at an energy point, where free parameters $m_{J_{t}}$ and $\lambda$ can be varied to minimize $\chi^{2}_{\rm Model}$.
+The fitting function $R_{b}^{\rm Model}(\mathcal{L}_{i},\sqrt{s_{i}})$ gives a theoretical $R_{b}$ value at an energy point, where free parameters $m_{J_{t}}$ and $\lambda$ can be varied to minimize $\chi^{2}_{\rm Model}$.
 […]
-Figure 1: The blue dashed line represents $R_{b}$ distribution in SM without $J_{t}$ model, while the red solid line represents $R_{b}$ distribution in SM with $J_{t}$.
-Additionally, the statistical significance of three energy points is indicated.
-The insert shows the LO Feynman diagram for the process $e^{+}e^{-}\rightarrow q\bar{q}$, with the corresponding propagator represented by a photon, a Z boson, or a $J_{t}$.
+Figure 1: The blue dashed line represents $R_{b}$ distribution in SM without $J_{t}$, the red solid line represents $R_{b}$ distribution in SM with $J_{t}$, and the statistical significance of the three energy points is indicated.
+The upper left insert shows $R_{b}$ values for $m_{t}^{\rm Input}$ and $m_{t}^{\rm Input}\pm 30~{}\text{MeV}$ near the second energy point.
+The lower right insert shows the LO Feynman diagram for $e^{+}e^{-}\rightarrow q\bar{q}$, where the propagator can be a photon, a $Z$, or a $J_{t}$.
 Figure 2: Significance of the $J_{t}$ signal varies with $m_{J_{t}}$ and $\lambda$ within the $1\sigma$ range.
 The main sources of systematic uncertainties in the $R_{b}$ measurements are the energy scale of $\sqrt{s}$, energy spread, detection efficiency, and backgrounds.
-All systematic uncertainties are more than an order of magnitude smaller than the statistical uncertainties (Supplementary Material), leading to a $J_{t}$ signal significance of $6.25\sigma$ even when considering the total systematic uncertainties.
+All systematic uncertainties are more than an order of magnitude smaller than the statistical uncertainties (End Matter), leading to a $J_{t}$ signal significance of $6.25\sigma$ even when the total systematic uncertainties are taken into account.
 When the true values of $m_{J_{t}}$ and $\lambda$ vary within their respective $1\sigma$ ranges, we determine the corresponding significance of the $J_{t}$ signal.
@@
 2.
-A significance level greater than $5\sigma$ indicates that the probability of SM with $J_{t}$ is incorrect compared to SM without $J_{t}$ being correct is less than $5.7\times 10^{-7}$.
-Extracting the top quark mass $m_{t}^{\text{1S}}$
-The top quark mass $m_{t}^{\text{1S}}$ (along with the Coulombic coefficient $\lambda$) is extracted by minimizing $\chi^{2}_{\text{SM with}~{}J_{t}}$.
+Extracting $m_{t}^{\text{1S}}$—The top quark mass $m_{t}^{\text{1S}}$ (along with the coefficient $\lambda$) is extracted by minimizing $\chi^{2}_{\text{SM with}~{}J_{t}}$.
 The statistical uncertainty $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$ is found to be $\rm 32.2~{}MeV$, primarily determined by the second energy point $\sqrt{s_{2}}$.
-The systematic uncertainties, $({\Delta m_{t}^{\text{1S}}})_{\rm syst.}$, arise from several sources, including the energy scale and energy spread, detection efficiency, backgrounds, SM parameters, higher order corrections, and others (luminosity measurements and backgrounds from two-photon and $WWZ$ processes).
+The systematic uncertainties, $({\Delta m_{t}^{\text{1S}}})_{\rm syst.}$, arise from several sources, including the energy scale and energy spread, detection efficiency, SM parameters, higher order corrections, and others (luminosity measurements and backgrounds from two-photon and $WWZ$ processes).
 The uncertainty of $\sqrt{s_{2}}$ energy scale is estimated to be 2.0 MeV at $2m_{t}^{\text{1S}}$ [50], which directly enters into the $m_{J_{t}}$ and contributes 1.0 MeV to $\Delta m_{t}^{\text{1S}}$.
 The center of mass energy spread is 458.4 MeV at $\sqrt{s_{2}}$ [16, 14], with an estimated uncertainty of 10% [16, 50], contributing an additional 0.2 MeV to ${\Delta m_{t}^{\text{1S}}}$.
-Experimentally, the background, as well as the detection efficiency and the weak mixing angle $\theta_{\rm W}$, can be determined primarily by the data from $Z$, $WW$, and $HZ$ peaks directly translated to the $m_{J_{t}}$ region [46].
-The detection efficiency related uncertainty in $R_{b}$ is ${\varepsilon_{\text{hadron}}^{udsc}}/{\varepsilon_{\text{tag}}^{b}}$ (Supplemental material), which is consistent that in the $R_{b}$ measurement at the Z pole.
-After including uncertainties from Monte Carlo (MC) statistics and energy-momentum requirement, the uncertainty associated with the detection efficiency in $R_{b}$ is $19.1\times 10^{-6}$ (Supplemental material).
-This is about half of the total uncertainty of the $R_{b}$ measurement at the Z-pole at CEPC, which is $4\times 10^{-5}$ [12].
+Experimentally, the backgrounds, as well as the detection efficiency and the weak mixing angle $\theta_{\rm W}$, can be determined primarily by the data from $Z$, $WW$, and $HZ$ peaks, which are directly translated to the $m_{J_{t}}$ region [51].
+The detection efficiency related uncertainty in $R_{b}$ is ${\varepsilon_{\text{hadron}}^{udsc}}/{\varepsilon_{\text{tag}}^{b}}$, which is consistent that in the $R_{b}$ measurement at the $Z$ pole.
+After including uncertainties from Monte Carlo statistics and energy-momentum requirement, the uncertainty associated with the detection efficiency in $R_{b}$ is $19.1\times 10^{-6}$ (End Matter).
+This is about half of the total uncertainty of the $R_{b}$ measurement at the $Z$ pole at CEPC, which is $4\times 10^{-5}$ [13].
 Consequently, this contributes 1.3 MeV to $\Delta m_{t}^{\text{1S}}$.
@@
 The NNLO correction of $R_{b}$ yields a contribution of 0.1 MeV to $\Delta m_{t}^{\text{1S}}$.
-The contribution of other sources, such as luminosity and backgrounds from two-photon and $WWZ$ processes, is less than 0.2 MeV for $\Delta m_{t}^{\text{1S}}$ (Supplemental material).
+The contribution of other sources, such as luminosity and backgrounds from two-photon and $WWZ$ processes, is less than 0.2 MeV for $\Delta m_{t}^{\text{1S}}$ (End Matter).
 Assuming these sources are independent, we combine them in quadrature, yielding a total systematic uncertainty of $({\Delta m_{t}^{\text{1S}}})_{\rm syst.}={1.9}$ MeV, as summarized in Table 1.
@@
 [17, 16] are six-body systems from $e^{+}e^{-}\to t\bar{t}\to b\bar{b}WW$, where $W$ decays into quark pairs or lepton-neutrino pairs, while our method focuses on two-body final states $b\bar{b}$.
-Table 1: Systematic uncertainties of the $m_{t}$ measurement in MeV. The “Other” category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
+Table 1: Systematic uncertainties of the $m_{t}$ measurement in MeV. The “Other" category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
 $\Delta m_{t}$ (MeV) FCC-ee [17] CEPC [16] This work
@@
 Background $10-20$ $4-18$ 0.9
-SM parameters ($\theta_{\rm W},~{}\alpha_{s}$) 3.2 17 $0.4$
+SM parameters ($\theta_{\rm W},\alpha_{s}$) 3.2 17 0.4
 Higher order corrections $40-45$ $9-26$ 0.1
@@
 Total syst. $40-74$ $23-58$ 1.9
-The upcoming $e^{+}e^{-}$ colliders CEPC and FCC-ee aim to achieve integrated luminosities of 1000 fb-1 [13] and 1700 fb-1 [14], respectively, near the $t\bar{t}$ production threshold.
+The upcoming $e^{+}e^{-}$ colliders aim to achieve integrated luminosities of 1,000 fb-1 at CEPC [14] and 1,700 fb-1 at FCC-ee [15], respectively, near the $t\bar{t}$ production threshold.
 With these luminosities, the systematic uncertainty on $m_{t}^{\text{1S}}$ is expected to remain at 1.9 MeV for both CEPC and FCC-ee.
-If the luminosities are chosen as $\mathcal{L}=(80,~{}900,~{}20)~{}{\rm fb}^{-1}$ and $\mathcal{L}=(80,~{}1600,~{}20)~{}{\rm fb}^{-1}$ for CEPC and FCC-ee at (340.076, 341.598, 342.771) GeV, respectively, our results show that, the statistical uncertainty, $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$, is estimated to be 18.6 MeV for CEPC and 13.9 MeV for FCC-ee, or 11.2 MeV when data from both colliders are combined.
-A detailed comparison of the top quark mass measurement methods at CEPC [15], FCC-ee [16], and this work is presented in Table 2.
-Table 2: Detailed comparison of the top quark mass measurement methods: final states, signal and background cross sections, and uncertainty estimation at CEPC [15], FCC-ee [16], and this work.
+If the luminosities are chosen as $\mathcal{L}=(80,900,20)~{}{\rm fb}^{-1}$ and $\mathcal{L}=(80,1,\!600,20)~{}{\rm fb}^{-1}$ for CEPC and FCC-ee at (340.076, 341.598, 342.771) GeV, respectively, our results show that, the statistical uncertainty, $({\Delta m_{t}^{\text{1S}}})_{\rm stat.}$, is estimated to be 18.6 MeV for CEPC and 13.9 MeV for FCC-ee, or 11.2 MeV when data from both colliders are combined.
+A detailed comparison of the top quark mass measurement methods at CEPC [16], FCC-ee [17], and in this work is presented in Table 2.
+Table 2: Detailed comparison of the top quark mass measurement methods: final states, signal and background cross sections, and uncertainty estimation at CEPC [16], FCC-ee [17], and in this work.
 Method CEPC [16] FCC-ee [17] This work
@@
 Measured quantity $\sigma({t\bar{t}})$ $\sigma({t\bar{t}})$ $R_{b}=\sigma({b\bar{b}})/\sigma({q\bar{q}})$
-$\sigma_{\rm signal}$ $\sim 450~{}$fb [17] $\sim 450~{}$fb [17]   $\sim 900$ fb ($b\bar{b}$), $\sim$ 6000 fb ($q\bar{q}$)
-[…]
+$\sigma_{\rm signal}$ $\sim 450~{}$fb [18] $\sim 450~{}$fb [18]   $\sim 900$ fb ($b\bar{b}$), $\sim$ 6,000 fb ($q\bar{q}$)
+Signal over background $\sim 2.9$ [18] $\sim 2.9$ [18] $\sim 89$, $\sim 12$
 $\sqrt{s}$ (GeV) $\sim 343~{}$GeV $\sim 343~{}$GeV $\sim 341~{}$GeV
-Integrated luminosity 100 fb-1 200 fb-1 $400-2700$ fb-1
+Integrated luminosity 100 fb-1 200 fb-1 $400-2,\!700$ fb-1
 $(\Delta m_{t})_{\rm stat.}$ $9$ MeV $9$ MeV $11.2-32.2$ MeV
@@
 $(\Delta m_{t})_{\rm total}$     $25-59$ MeV     $45-75$ MeV $11.3-32.3$ MeV
-Discussion and Outlook
-In this study, we propose a method for both discovering toponium and precisely measuring the top quark mass $m_{t}^{\text{1S}}$ using the observable $R_{b}$ near the energy threshold $2m_{t}^{\text{1S}}$ in $e^{+}e^{-}$ collisions.
-The significance of the toponium signal exceeds $5\sigma$, including both statistical and systematic uncertainties with the integrated luminosity of 400 $\rm fb^{-1}$ at around 341 GeV.
+Discussion and Outlook—In this Letter, we propose a method for both discovering toponium and precisely measuring the top quark mass $m_{t}^{\text{1S}}$ using the observable $R_{b}$ near the energy threshold $2m_{t}^{\text{1S}}$ in $e^{+}e^{-}$ collisions.
+The significance of the toponium signal exceeds $5\sigma$, including both statistical and systematic uncertainties with the integrated luminosity of 400 $\rm fb^{-1}$ around 341 GeV.
 The $m_{t}^{\text{1S}}$ can be measured with a total uncertainty of only 32.3 MeV, including a systematic uncertainty of 1.9 MeV.
 This enables a top quark mass measurement with an uncertainty reduced by a factor of 10 compared to the current precision [2].
-Moreover, our method improves the systematic uncertainty by more than a factor of 150 compared to the most precise current measurements [2], and by at least a factor of 12 compared to other existing methods [15, 16].
-As the smallest quantum bound state and simplest hadron, toponium provides a unique opportunity to probe fundamental principles of quantum mechanics [3, 5], including particle-wave duality, entanglement, and spin dynamics, while offering novel insights into the nature of confinement in QCD.
-Additionally, it could play a critical role in influencing the cooling rate and supercooled electroweak phase transition in the early universe [31, 32, 33].
+Additionally, polarized beams or the well separation of $b$-jet and $\bar{b}$-jet in detection can reduce the integrated luminosities by an order of magnitude.
+As the smallest quantum bound state and simplest hadron, toponium provides a unique opportunity to probe fundamental principles of quantum mechanics [3, 5], including particle-wave duality, entanglement, and spin dynamics, while offering novel insights into the nature of confinement in QCD [28].
 A precise measurement of $m_{t}$ is crucial in particle physics, as it directly impacts the stability of SM vacuum and determines whether the universe exists in a metastable state [11, 12].
-However, it should be noted that while our method offers a high precision, it is sensitive to the parameter $\lambda$, which governs the coupling between $J_{t}$ and both $b\bar{b}$ and $e^{+}e^{-}$.
-A precise understanding of $\lambda$ is essential, and can be achieved by lattice QCD or higher order PQCD calculations.
-Additionally, polarized beams or the well separation of $b$-jet and $\bar{b}$-jet in detection can reduce the integrated luminosities by an order of magnitude while maintaining the accuracy of the $m_{t}$ measurements.
-Alternatively, they can increase the accuracy of the $m_{t}$ measurements by a factor of three without altering the integrated luminosities.
-Furthermore, comparisons with the multi-body final state approach to top quark mass measurements could help to validate and complement the results obtained here.
-Ultimately, the goal is to integrate this method into a broader framework for discovering new bound states and measuring the masses of fundamental particles, providing a robust and comprehensive understanding of quantum mechanics and particle physics.
-Acknowledgements.
-We thank Profs.
+Acknowledgments—We thank Profs.
 Kuang-Ta Chao, Gang Li, Yan-Qing Ma, Manqi Ruan, Xiaohu Sun, and Shi-Lin Zhu for valuable and helpful discussions.
 This work is supported in part by National Key Research and Development Program of China under Contract No.
-2020YFA0406300, and No.
 2024YFA1610503, National Natural Science Foundation of China (NSFC) under contract No.
-11975076, No.
 12161141008, No.
-12135005, No.
-12075018, and No.
+12135005, and No.
 12405102, the Fundamental Research Funds for the Central Universities Grant No.
@@
 2024M750049.
-Appendix A Supplemental material
-A.1 Top quark and toponium
-The most precise $m_{t}^{\text{OS}}$ measurement method is the top quark pair production at the $2m_{t}^{\text{OS}}$ threshold at the future $e^{+}e^{-}$ colliders [15, 47, 48, 49, 50, 51, 52, 17, 53, 54, 55], such as CEPC [12], Compact Linear Collider [56], FCC-ee [14], and International Linear Collider [57].
-The cross sections of $e^{+}e^{-}\to t\bar{t}$ near the threshold are calculated to NNLO [53, 58, 59, 60] and NNNLO [61, 54, 62, 63, 61].
-[…]
-A.2 Theoretical calculation
-A.2.1 Modeling $J_{t}$ interaction
-The ${J_{t}}$ bound state results from the Schrödinger equation with a potential between $t$ and $\bar{t}$, which is simplified to a Coulomb plus linear potential based on lattice QCD results [39, 40]:
-$V(r)=-\frac{\lambda}{r}+\sigma r,$ (10)
-where $r$ is the distance between $t$ and $\bar{t}$, the Coulombic coefficient is $\lambda=0.285\pm 0.011$, and the string tension is $\sigma=0.206~{}{\rm GeV}^{2}$ [39, 40].
-The Bohr radius of toponium is given by $r_{B}={2}/{(m_{t}^{\text{OS}}\lambda)}$, which evaluates to ${1}/(24.584~{}\rm{GeV})$.
-The $\sigma r$ term in the potential can be neglected, as its contribution to the binding energy of $J_{t}$ is $-\langle\sigma r\rangle=-{3\sigma r_{B}}/{2}=-0.013~{}{\rm GeV}$.
-Additionally, the simplest bosonic string model predicts the same potential as given in Eq.(10), with $\lambda=\pi/12=0.262$ [100].
-This prediction is confirmed by result from SU(3) lattice gauge theory, which suggests a slightly higher value of $\lambda\sim 0.293$ [101].
-The binding energy $B_{J_{t}}$ can also be calculated in perturbative QCD (PQCD), yielding $(1.57_{\rm LO}+0.65_{\rm NLO}+0.49_{\rm NNLO}+0.25_{\rm NNNLO})~{}{\rm GeV}=2.96$ GeV at NNNLO with a renormalization scale of $\mu=32.6$ GeV for $m_{J_{t}}$ and $\mu=30$ GeV for $m_{t}^{\text{OS}}$ [102].
+Appendix A End Matter
+Theoretical calculation—The binding energy $B_{J_{t}}$ can also be calculated in perturbative QCD (PQCD), yielding $(1.57_{\rm LO}+0.65_{\rm NLO}+0.49_{\rm NNLO}+0.25_{\rm NNNLO})~{}{\rm GeV}=2.96$ GeV at NNNLO with a renormalization scale of $\mu=32.6$ GeV for $m_{J_{t}}$ and $\mu=30$ GeV for $m_{t}^{\text{OS}}$ [52].
 The difference between $m_{t}^{\text{OS}}$ and $m_{t}^{\text{1S}}$ is calculated to be $(0.428_{\rm LO}+0.368_{\rm NLO}+0.262_{\rm NNLO}+(0.174\pm 0.0004)_{\rm NNNLO})$ GeV $=(1.232\pm 0.0004)$ GeV at NNNLO with $\mu=163.508$ GeV [25].
 With the center value of $m_{t}^{\text{OS}}=172.52~{}{\rm GeV}$ [2], the $\tt{QQbar\_threshold}$ package [53, 54, 55] yields:
-$B_{J_{t}}^{\rm PQCD}=(1.680_{\rm LO}+0.657_{\rm NLO}+0.504_{\rm NNLO}+0.243_{\rm NNNLO})~{}{\rm GeV}=3.084~{}{\rm GeV}$ (11)
-at NNNLO with $\mu=1/r_{B}=24.584$ GeV.
-The NkLO term is at the order of $(\alpha_{s}(\mu)/\pi)^{k}\sim 0.04^{k}$ for $k=0,1,2,3,...$.
+$B_{J_{t}}^{\rm PQCD}=$ $(1.680_{\rm LO}+0.657_{\rm NLO}+0.504_{\rm NNLO}$ (10)
+$+0.243_{\rm NNNLO})~{}{\rm GeV}=3.084~{}{\rm GeV}$
+at NNNLO with $\mu=1/r_{B}=24.584$ GeV, where $r_{B}$ is the Bohr radius.
 However, the convergence of the perturbative series up to NNNLO is still unsatisfactory, necessitating consideration of higher order corrections.
-Including contributions from the Higgs and quantum electrodynamics (QED), the binding energy is shifted by $0.110$ GeV [104], thereby bringing the NNNLO binding energy in Eq.(11) to 3.194 GeV.
-Such result can be compared to that of $(3.503\pm 0.270)$ GeV in Eq.(Modeling $J_{t}$ Interaction).
-We introduce the Lagrangian density $\mathcal{L}_{J_{t}}$ of the field $J_{t}$ into that of SM, $\mathcal{L}_{\text{SM}}$ [36]:
-$\mathcal{L}_{J_{t}}$ $=$ $\sum_{f=b,e}\bar{f}\left(g^{V}_{{J_{t}}ff}\gamma_{\mu}+g^{A}_{{J_{t}}ff}\gamma_{\mu}\gamma^{5}\right)f{J_{t}}^{\mu}+\ldots,$ (12)
-The ellipse represents the kinetic energy and mass term of $J_{t}$, as well as the other interactions between $J_{t}$ and the fundamental particles.
-The factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$ can be determined with non-relativistic QCD (NRQCD)  [38]:
-$g^{A}_{{J_{t}}ee}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\frac{2\pi\alpha m_{t}^{2}\left(3\csc^{2}\theta_{\rm W}-8\right)\sec^{2}\theta_{\rm W}}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)},$
-$g^{V}_{{J_{t}}ee}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\frac{8\pi\alpha\left(m_{t}^{2}\left(6\cos\left(2\theta_{\rm W}\right)-9\right)\csc^{2}\left(2\theta_{\rm W}\right)+2m_{Z}^{2}\right)}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)},$
-$g^{A}_{{J_{t}}bb}=\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\left(\frac{2\pi\alpha m_{t}^{2}\left(m_{t}^{2}+2m_{W}^{2}\right)\csc^{2}\theta_{\rm W}}{\sqrt{3}m_{W}^{2}\left(m_{t}^{2}+m_{W}^{2}\right)}+\frac{8\pi\alpha m_{t}^{2}\left(4\cos\left(2\theta_{\rm W}\right)-1\right)\csc^{2}\left(2\theta_{\rm W}\right)}{\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)}\right),$
-$g^{V}_{{J_{t}}bb}=-\frac{\sqrt{\lambda^{3}}}{8\sqrt{\pi}}\left(\frac{16\pi\alpha}{3\sqrt{3}}+\frac{2\pi\alpha m_{t}^{2}\left(m_{t}^{2}+2m_{W}^{2}\right)\csc^{2}\theta_{\rm W}}{\sqrt{3}m_{W}^{2}\left(m_{t}^{2}+m_{W}^{2}\right)}\right.$
-$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\vspace{4cm}\left.+\frac{8\pi\alpha m_{t}^{2}\left(2\cos\left(2\theta_{\rm W}\right)+4\cos\left(4\theta_{\rm W}\right)+3\right)\csc^{2}\left(2\theta_{\rm W}\right)}{3\sqrt{3}\left(4m_{t}^{2}-m_{Z}^{2}\right)}\right).$ (13)
-Here $\alpha$ is the running QED coupling constant.
-The Lagrangian density $\mathcal{L}_{J_{t}}$ is encoded in $\tt FeynRules$ [105].
-The calculations can be done with the following packages: $\tt FeynArts$ [106], $\tt FeynCalc$ [107], $\tt MadGraph$ [108], and $\tt Whizard$ [109].
-The width $\Gamma_{J_{t}}$ of $J_{t}$, dominated by twice the width of the top quark $t$, is:
-$\Gamma_{J_{t}}$ $=$ $2\Gamma_{t}^{\text{OS}}\left(1-\frac{\lambda^{2}}{8}\right)+\Gamma_{\rm Annihilation}=(2.603\pm 0.019)~{}{\rm GeV},$ (14)
-where $\Gamma_{t}^{\text{OS}}=(1.3148^{+0.003}_{-0.005}+0.027(m_{t}^{\text{OS}}-172.69))~{}{\rm GeV}$ [42, 110], and the coefficient $\left(1-\lambda^{2}/8\right)$ comes from corrections for the fact that bound top quarks are different from free top quarks  [111].
-In addition to the corrected width of top quark, we also consider the two-body annihilation decay width of $J_{t}$, $\Gamma_{\rm Annihilation}=0.0093~{}{\rm GeV}$, based on NRQCD [38].
-Since the branching fraction of $J_{t}$ to $b\bar{b}$ is relatively large, we choose the measured quantity $R_{b}$ defined in Eq.(1). The parameters are chosen as [112, 2, 113]:
-$m_{t}^{\text{OS}}=(172.52\pm 0.33)~{}{\rm GeV},$ (15)
-$m_{W}=(80.3692\pm 0.0133)~{}{\rm GeV},$
-$m_{Z}=(91.1880\pm 0.0020)~{}{\rm GeV},$
-$m_{H}=(125.20\pm 0.11)~{}{\rm GeV},$
-$m_{b}(m_{J_{t}})=(2.568\pm 0.010)~{}{\rm GeV},$
-$m_{c}(m_{J_{t}})=(0.575\pm 0.003)~{}{\rm GeV},$
-$\alpha(m_{J_{t}})=1/(126.04\pm 0.01),$
-$\alpha_{s}(m_{J_{t}})=0.09844\pm 0.00062,$
-$\cos\theta_{\rm W}=m_{W}/m_{Z}.$
-The $\alpha(m_{J_{t}})$, the running QCD coupling constant $\alpha_{s}(m_{J_{t}})$, the running charm quark mass $m_{c}(m_{J_{t}})$, and the running bottom quark mass $m_{b}(m_{J_{t}})$ are calculated from $\alpha(m_{Z})=1/(127.930\pm 0.008)$ and $\alpha_{s}(m_{Z})=0.1180\pm 0.0009$ [112] with $\tt{RunDec}$ package [114].
-The Cabibbo-Kobayashi-Maskawa (CKM) matrix is also to be employed.
-However, the offset of $V_{tb}$ from 1 is much smaller than the uncertainty of $\lambda$, and the contribution of $V_{cb}$ is suppressed to less than $1/100$ of the statistical uncertainty of the $R_{b}$.
-Therefore, the CKM matrix can be approximated as the unit matrix.
-Only the $R_{b}$ uncertainty originating from $\theta_{\rm W}$ needs to be preserved, as all others are smaller than $10^{-6}$.
-The production cross section of pseudoscalar toponium $\eta_{t}$ at LHC is $6.24$ pb at 13 TeV with the same mass and wave function as $J_{t}$ here and the parton distribution $\tt{NNPDF40\_lo\_as\_01180}$ [115], which is consistent with $(7.1\pm 0.8)$ pb as recently reported by CMS [41].
-This result underscores the significance of future $e^{+}e^{-}$ colliders, such as CEPC and FCC-ee, in probing toponium with an unprecedented precision.
-By leveraging the high luminosities and clean environments of these colliders, future experiments could significantly improve upon the detection and measurement of such states.
-A.2.2 Theoretical prediction of $R_{b}$ near $2m_{t}^{\rm 1S}$
-The experimental cross section, denoted by the symbol $\sigma$, is given by the following expression, as presented in Refs. [43, 63]:
+The running parameters are chosen at the energy scale of $m_{J_{t}}$ for both theoretical calculations and experimental simulations [56, 2, 57, 58]:
+$m_{b}=2.568(10)~{}{\rm GeV}$ , $~{}m_{c}=0.575(3)~{}{\rm GeV},$
+$m_{H}=125.20(11)~{}{\rm GeV}$ , $~{}\alpha_{s}=0.09844(62),$
+$m_{Z}=91.1880(20)~{}{\rm GeV}$ , $~{}\alpha=1/126.04(1),$
+$m_{W}=80.3692(133)~{}{\rm GeV}$ , $~{}\cos\theta_{\rm W}=m_{W}/m_{Z}.$ (11)
+The experimental cross section, denoted by the symbol $\sigma$, is given by the following expression [48, 54]:
 […]
-In this formula, $\delta_{\sqrt{s}}$ is the center of mass energy spread, which is proportional to the square of $\sqrt{s}$ [15, 13].
-Due to $\delta_{\sqrt{s}}$ is $460.30.186\%\times\sqrt{2}\times 175~{}{\rm GeV}=$ 460.3 MeV for FCC-ee at 350 GeV [14], and $0.20\%\times\sqrt{2}\times 180~{}$GeV$=509.1~{}$MeV for CEPC at 360 GeV [13], we can get $\delta_{\sqrt{s}}$ is about 440 MeV for FCC-ee and 460 MeV for CEPC at around $m_{J_{t}}$.
-The ISR factor $F(x,\sqrt{s^{\prime}})$ is given in Refs.
+×$\times\int dx~{}F(x,\sqrt{s^{\prime}})\bar{\sigma}(\sqrt{xs^{\prime}})=f_{\rm ISR}~{}\sigma_{\rm Born}.$ (12)
+In this formula, $\delta_{\sqrt{s}}$ is the center of mass energy spread, and the ISR factor $F(x,\sqrt{s^{\prime}})$ is given in Refs.
 [59, 54, 60].
 Then, we can obtain $f_{\rm ISR}$, the factor that includes the effects of energy spread and ISR.
-Figure 3: The LO Feynman diagrams of the $e^{+}e^{-}\rightarrow q\bar{q}$ process comprise three diagrams: the first and second diagrams correspond to SM, while the third diagram corresponds to the $J_{t}$ contribution.
-The LO Feynman diagrams of the $e^{+}e^{-}\rightarrow q\bar{q}$ are shown in Fig. 3. The LO amplitude of $e^{+}e^{-}\to q\bar{q}$ can be written as:
-${\mathcal{M}}(e^{+}e^{-}\to q\bar{q})$ $=$ $\frac{g^{VV}_{q}}{s}\bar{e}.\gamma^{\mu}.e~{}\bar{q}.\gamma_{\mu}.q+\frac{g^{VA}_{q}}{s}\bar{e}.\gamma^{\mu}.e~{}\bar{q}.\gamma_{\mu}.\gamma^{5}.q$ (17)
-$+~{}\frac{g^{AV}_{q}}{s}\bar{e}.\gamma^{\mu}.\gamma^{5}.e~{}\bar{q}.\gamma_{\mu}.q+\frac{g^{AA}_{q}}{s}\bar{e}.\gamma^{\mu}.\gamma^{5}.e~{}\bar{q}.\gamma_{\mu}.\gamma^{5}.q,$
-where the coefficient $g^{VA}_{q}$ (similar to $g^{VV}_{q}$, $g^{AV}_{q}$, and $g^{AA}_{q}$ ) is:
-$g^{VA}_{q}$ $=$ $\sum_{j=\gamma,~{}Z,~{}J_{t}}{g^{V}_{jee}}^{*}\times g^{A}_{jqq}\times\frac{s}{s-m_{j}^{2}+i\Gamma_{j}m_{j}}.$ (18)
-The ${g^{V}_{jee}}^{*}$ denotes the complex conjugate of $g^{V}_{jee}$.
-The factors $g^{V}_{{J_{t}}ff}$ and $g^{A}_{{J_{t}}ff}$, mass $m_{J_{t}}$, and the width $\Gamma_{J_{t}}$ are given in Subsec.
-A.2.1, and those for $\gamma$ and $Z$ are given in Ref.[112].
-Then the cross section of $\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})$ is  [118]:
-$\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})=\frac{1}{12\pi s}\left(f^{A}_{q}\left(\left|g_{q}^{AA}\right|^{2}+\left|g_{q}^{VA}\right|^{2}\right)+f^{V}_{q}\left(\left|g^{AV}_{q}\right|^{2}+\left|g_{q}^{VV}\right|^{2}\right)\right).$ (19)
-The factor $f^{A}_{q}$ ($f^{V}_{q}$) includes the correction factor of the axis vector current $\bar{b}.\gamma_{\mu}.\gamma^{5}.b$ (vector current $\bar{b}.\gamma_{\mu}.b$), which is 1 at LO.
-The factors $f_{A}$ and $f_{V}$ were calculated to NLO  [118], NNLO [119], and NNNLO [120, 61], and the results are described in Ref. [121]:
-$f^{A}_{q}$ $=$ $f^{0}_{q}+(\delta_{qc}+\delta_{qb})f^{A}_{q,2},$
-$f^{V}_{q}$ $=$ $f^{0}_{q}+(\delta_{qc}+\delta_{qb})f^{V}_{q,2},$
-$f^{0}_{q}$ $=$ $1+a_{s}+\frac{3}{4}Q_{q}^{2}\frac{\alpha(\sqrt{s})}{\pi}\left(1-\frac{a_{s}}{3}\right)+a_{s}^{2}\left[\frac{365}{24}-11\zeta_{3}+n_{f}\left(-\frac{11}{12}+\frac{2}{3}\zeta_{3}\right)\right]+{\mathcal{O}}(a_{s}^{3})$
-$=$ $1+a_{s}+\frac{3}{4}Q_{q}^{2}\frac{\alpha(\sqrt{s})}{\pi}\left(1-\frac{a_{s}}{3}\right)+a_{s}^{2}\left(1.9857-0.1153n_{f}\right)$
-$\hskip 28.45274pt+a_{s}^{3}\left(-6.6369-1.2001n_{f}-0.005178n_{f}^{2}\right)+{\mathcal{O}}\left(a_{s}^{3}\frac{m_{b}^{2}}{s},a_{s}^{4}\right),$
-$f^{A}_{q,2}$ $=$ $\frac{m_{q}^{2}}{s}\left[-6-22a_{s}+a_{s}^{2}\left(-\frac{8221}{24}+57\zeta_{2}+117\zeta_{3}+n_{f}\left(\frac{151}{12}-2\zeta_{2}-4\zeta_{3}\right)\right)+{\mathcal{O}}(a^{3}_{s})\right]$
-$=$ $\frac{m_{q}^{2}}{s}\left[-6-22a_{s}+a_{s}^{2}\left(-108.140+4.48524n_{f}\right)\right]+{\mathcal{O}}\left(\frac{m_{b}^{4}}{s^{2}},a_{s}^{3}\frac{m_{b}^{2}}{s}\right),$
-$f^{V}_{q,2}$ $=$ $\frac{m_{q}^{2}}{s}\left[12a_{s}+a_{s}^{2}\left(\frac{253}{2}-\frac{13}{3}n_{f}\right)+{\mathcal{O}}(a^{3}_{s})\right]$ (20)
-$=$ $\frac{m_{q}^{2}}{s}\left[12a_{s}+a_{s}^{2}\left(126.500-4.33333n_{f}\right)\right]+{\mathcal{O}}\left(\frac{m_{q}^{4}}{s^{2}},a_{s}^{3}\frac{m_{q}^{2}}{s}\right),$
-where the $\delta_{qb}$ ($\delta_{qc}$) indicates the need to include the bottom quark mass $m_{b}$ (the charm quark mass $m_{c}$) and $m_{u}=m_{d}=m_{s}=0$, $a_{s}=\alpha_{s}(\sqrt{s})/\pi$, $Q_{b}=Q_{s}=Q_{d}=-1/3$, and $Q_{u}=Q_{c}=Q_{t}=-1/3$.
-The $\zeta$ is Riemann Zeta function, and $n_{f}=5$ is the flavor number of active quarks.
-The terms $(\alpha_{s}/\pi)^{n}\times(m_{b}^{2}/s)^{0}$ with $n>3$ are canceled out in $R_{b}$ defined in Eq.(1), so the contribution of higher order corrections in $R_{b}$ is the ${m_{b}^{4}}/{s^{2}}\sim(\alpha_{s}/\pi)^{3}\times(m_{b}^{2}/s)<10^{-7}$.
-Due to $\sigma_{\rm Born}(e^{+}e^{-}\to ggg)<10^{-2}~{}$fb is very small, we dropped it from $\sigma_{\rm Born}(e^{+}e^{-}\to q\bar{q})$ in Eq.(1).
-In SM, the LO, NLO, NNLO, and NNNLO predictions of $R_{b}$ at $\sqrt{s}$ around $m_{J_{t}}$ are:
-$R_{b,~{}{\rm LO}}^{\rm SM}$ $=$ $\left(151550.209-27.365\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
-$R_{b,~{}{\rm NLO}}^{\rm SM}$ $=$ $\left(151499.646-27.362\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
-$R_{b,~{}{\rm NNLO}}^{\rm SM}$ $=$ $\left(151500.422-27.363\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6},$
-$R_{b,~{}{\rm NNNLO}}^{\rm SM}$ $=$ $\left(151500.457-27.364\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}+0.127\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\right)\times 10^{-6}.$ (21)
-Considering that future $R_{b}$ measurement at $365$ GeV with $1.5~{}{\rm ab}^{-1}$ data at FCC-ee will have a precision of $1.4\times 10^{-4}$ [35], the NLO calculation of $R_{b}$ is sufficient, and the higher order QCD corrections of $R_{b}$ are much smaller than the systematic uncertainties and can be safely ignored.
-The NLO Born cross sections of $e^{+}e^{-}\to b\bar{b}$ in SM without and with $J_{t}$ are shown in Fig.
-4.
-Figure 4: The NLO Born cross sections of $e^{+}e^{-}\to b\bar{b}$ in SM without and with $J_{t}$.
-A.2.3 Measuring $R_{b}$ near $2m_{t}^{\rm 1S}$
+In SM, the LO [56], NLO [61], NNLO [62] and NNNLO [63, 64, 65] predictions of $R_{b}$ at $\sqrt{s}$ around $m_{J_{t}}$ are:
+$R_{b}^{\rm SM}$ $=$ $\bigg{[}151,\!550.2_{\rm LO}-50.6_{\rm NLO}+0.8_{\rm NNLO}$ (13)
+$+0.0_{\rm NNNLO}-27.4\times\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}$
+$+0.1\times\left(\frac{\sqrt{s}-m_{J_{t}}}{\rm GeV}\right)^{2}+...\bigg{]}\times 10^{-6}.$
 The $R_{b}(\sqrt{s_{i}})$ is determined by the cross sections of $\sigma(e^{+}e^{-}\to b\bar{b})$ and $\sigma(e^{+}e^{-}\to q\bar{q})$ at the selected $\sqrt{s_{i}}$:
-$N_{\text{tag}}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q}\varepsilon_{\text{tag}}^{q}\sigma(e^{+}e^{-}\to q\bar{q}),$
-$N_{\rm hadron}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q}\varepsilon^{q}_{\rm hadron}\sigma(e^{+}e^{-}\to q\bar{q}).$ (22)
-The $N_{\text{tag}}$, $N_{\rm hadron}$, $\varepsilon^{q}_{\rm tag}$, and $\varepsilon^{q}_{\rm hadron}$ are defined after Eq.(7).
-Then the $R_{b}$ is defined in Eq.(7).
-If we ignore the small values $\varepsilon_{\text{hadron}}^{udsc}-\varepsilon_{\text{hadron}}^{b}$, $\varepsilon_{\text{tag}}^{udsc}$, and $1-{f_{\rm ISR}^{b}}/{f_{\rm ISR}^{udsc}}$ , we can get:
-$R_{b}\sim\frac{N_{\text{tag}}\varepsilon_{\text{hadron}}^{udsc}}{N_{\text{hadron}}\varepsilon_{\text{tag}}^{b}}.$ (23)
-This equation is used to estimate only the $R_{b}$ uncertainty associated with the detection efficiency.
-A.3 Experimental simulation
-In this study, MC simulations of $e^{+}e^{-}$ collisions have been performed.
-The physics events involving the physics process with or without $J_{t}$ are generated using $\tt Whizard$ [109], where the effects of ISR and energy spread have been taken into account.
-Subsequently, the generated parton-level events from $\tt Whizard$ are passed to $\tt Pythia6$ [122] for parton showering, hadronization, and the decay processes of the short-life hadrons, ensuring a realistic simulation of the final state particles.
-Finally, $\tt Delphes$ [123], configured with the CEPC detector conditions [12], is used to simulate the detector response, allowing for the reconstruction of observable quantities from the theoretical predictions.
-A.3.1 Event generation in pseudo-experiments
-The best option to identify $J_{t}$ is in the final state with double $b$-jets.
-However, jets originating from other quark flavors can sometimes be misidentified as $b$-jets [124] with the $b$-tagging algorithms [125], thus the contribution of other final states ($u\bar{u}$,${}d\bar{d}$$s\bar{s}$$c\bar{c}$) should be also generated.
-Moreover, due to limitations in detector performance, some jets with low energy or those produced at extreme angles relative to the detector’s central axis [12, 14] may go undetected, which results in an underestimation of the jet multiplicity in some events.
+$N_{\text{tag}}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q=udscb}\varepsilon_{\text{tag}}^{q}\sigma(e^{+}e^{-}\to q\bar{q}),$
+$N_{\rm hadron}(\sqrt{s_{i}})$ $=$ $\mathcal{L}_{i}\sum_{q=udscb}\varepsilon^{q}_{\rm hadron}\sigma(e^{+}e^{-}\to q\bar{q}).$ (14)
+The $N_{\text{tag}}$, $N_{\rm hadron}$, $\varepsilon^{q}_{\rm tag}$, and $\varepsilon^{q}_{\rm hadron}$ are defined after Eq.(7). Then the $R_{b}$ is defined in Eq.(7).
+Experimental simulation—In this Letter, MC simulations of $e^{+}e^{-}$ collisions have been performed.
+The physics events are generated using $\tt Whizard$ [66], where the effects of ISR and energy spread have been taken into account.
+Subsequently, $\tt Pythia6$ [67] is applied for parton showering, hadronization, and the decay processes of the short-life hadrons.
+Finally, $\tt Delphes$ [68], configured with the CEPC detector conditions [13], is used to simulate the detector response.
+To measure $R_{b}$, samples with 2 fermions ($u\bar{u},d\bar{d},s\bar{s},c\bar{c}$,$b\bar{b}$) are generated.
+Moreover, due to limitations in detector performance, some jets with low energy or those produced at extreme angles relative to the detector’s central axis [13, 15] may go undetected.
 To address this, several multi-jet backgrounds from $WW$, $ZZ$, and $ZH$ decays to four-quark final states are simulated.
 In order to enhance the sensitivity for $J_{t}$ signal, the event selection criteria are optimized as discussed below.
-A.3.2 Event selection criteria
 Jets from $q\bar{q}$ are clustered from all the well reconstructed tracks with an anti-kt algorithm [69].
@@
 Thus, the effective invariant mass of the final states $\sqrt{s^{\prime}}$ needs to be considered.
-Consequently, $\sqrt{s^{\prime}}$ in Eq.(16) can not be obtained directly by the energies of jets [44], but can be measured well by using the information of jet angular detection with an accuracy of $0.01\%$ [127].
-Based on the conservation of four-momentum of the initial and final states, $\sqrt{s^{\prime}}$ of the three-jet final states can be determined as:
-$\sqrt{s^{\prime}}=\begin{cases}\sqrt{s}-\sqrt{s}\frac{\cos\theta_{1}\sin\phi_{23}+\sin\theta_{1}\left(\cot\theta_{2}\sin\phi_{31}-\cot\theta_{3}\sin\phi_{21}\right)}{\left(\cos\theta_{1}+1\right)\sin\phi_{23}+\sin\theta_{1}\left(\cot\frac{\theta_{2}}{2}\sin\phi_{31}-\cot\frac{\theta_{3}}{2}\sin\phi_{21}\right)}&\text{(Forward)}\\ \sqrt{s}-\sqrt{s}\frac{\cos\theta_{1}\sin\phi_{23}+\sin\theta_{1}\left(\cot\theta_{2}\sin\phi_{31}-\cot\theta_{3}\sin\phi_{21}\right)}{\left(\cos\theta_{1}-1\right)\sin\phi_{23}+\sin\theta_{1}\left(\tan\frac{\theta_{3}}{2}\sin\phi_{21}-\tan\frac{\theta_{2}}{2}\sin\phi_{31}\right)}&\text{(Backward)}\end{cases},$ (24)
-where $\theta_{i}$ and $\phi_{i}$ are the polar and azimuthal angles of the $i-$th jet for $i=1,~{}2,~{}3$, and $\phi_{ij}=\phi_{i}-\phi_{j}$.
-Forward (backward) indicates that the momentum direction of the ISR photon or missing jet is the same as (opposite to) the direction of the electron beam.
-In the two cases of forward and backward directions, $1-{\sqrt{s^{\prime}}}/{\sqrt{s}}$ is positive in one case and negative in the other, where the positive value represents the physical solution.
-Similarly, in the two-jet final states, we can get:
-$s^{\prime}=\frac{\sin{\theta_{1}}+\sin{\theta_{2}}-|\sin{(\theta_{1}+\theta_{2})}|}{\sin{\theta_{1}}+\sin{\theta_{2}}+|\sin{(\theta_{1}+\theta_{2})}|}\times s,$ (25)
-which is consistent with the Eq.(1) of Ref. [44]. We require $\sqrt{s^{\prime}}>\sqrt{s}-10~{}{\rm GeV}$ here [128, 44, 63].
+We require $\sqrt{s^{\prime}}>\sqrt{s}-10~{}{\rm GeV}$ here [70, 49, 54].
 For the measurement of $R_{b}$, only events with at least two $b$-tagged jets [71] are categorized as $b\bar{b}$ events, while the remaining events are considered as light quark pair processes.
@@
 Events passing all of the criteria are recognized as signal events, and the detection efficiencies of different processes in the three energy points are shown in Table 3.
-Table 3: The experimental cross sections $\sigma(\sqrt{s})$ in fb, double $b$ tagged efficiencies $\varepsilon_{\text{tag}}(\%)$, and detection efficiencies $\varepsilon_{\text{hadron}}(\%)$ for different processes in $R_{b}$ measurements.
-Here $jj$ is $u\bar{u}$,${}d\bar{d}$$s\bar{s}$$c\bar{c}$, or $gg$.
+Systematic uncertainties—The systematic uncertainties, affecting the accuracy of $R_{b}$ measurements, arise from accelerator effects (e.g., $\sqrt{s}$ energy scale and energy spread), detector efficiencies (e.g., MC statistics, $\sqrt{s^{\prime}}$ and $p_{t}$ requirements), SM parameters ($\theta_{\rm W}$, $\alpha_{s}$, etc.), theoretical considerations (e.g., higher order corrections and ISR), and others (e.g., luminosity measurements and backgrounds from two-photon and $WWZ$ processes).
+Taking the $R_{b}$ measurement at the second energy point as an example, systematic uncertainties are discussed below.
+The uncertainty in the energy scale of $\sqrt{s_{2}}$ is estimated to be 2 MeV [50], contributing $14.6\times 10^{-6}$ to $\Delta R_{b}$.
+The energy spread at the second energy point is 458.4 MeV for CEPC, with a 10% uncertainty [16, 50] contributing an additional $2.7\times 10^{-6}$ to $\Delta R_{b}$.
+By increasing the size of simulated events, the statistical error can be reduced to less than $10^{-4}$, resulting in a contribution of $15.4\times 10^{-6}$ to $\Delta R_{b}$.
+The precisions of $p_{t}$ and $\sqrt{s^{\prime}}$ are controlled to within 0.10 GeV and 0.34 GeV [49, 72], respectively, leading to $\Delta R_{b}$ values of $9.1\times 10^{-6}$ and $6.7\times 10^{-6}$.
+The $WW$, $ZZ$, and $ZH$ related processes are well understood both theoretically and experimentally, with approximately 95% of such events rejected by selection criteria.
+The relative uncertainty in these backgrounds is less than $(\alpha_{s}/\pi)^{2}\sim 0.001$, contributing $13.9\times 10^{-6}$ to $\Delta R_{b}$.
+The uncertainty in $\sin^{2}\theta_{\rm W}$, achievable at $1\times 10^{-5}$ at CEPC [13, 50], contributes $6.0\times 10^{-6}$ to $\Delta R_{b}$.
+The uncertainty of $m_{b}$ in Eq.(A) contributes $0.2\times 10^{-6}$ to $\Delta R_{b}$, while the uncertainty from $\alpha_{s}$ is included in higher-order corrections.
+As shown in Eq.(13), the NNLO corrections increase $R_{b}$ by $1.2\times 10^{-6}$.
+The other sources include luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
+Their contributions to ${\Delta R_{b}}$ are small, with a combined estimation of less than $3.0\times 10^{-6}$.
+All the systematic uncertainties are summarized in Table 4.
+Table 3: The experimental cross sections $\sigma(\sqrt{s})$ (in fb), along with the detection efficiencies $\varepsilon_{\text{tag}}$ (in %, for double $b$-tagged events) and $\varepsilon_{\text{hadron}}$ (in %, for hadronic events), are provided for various processes in the $R_{b}$ measurements.
+Here, $jj$ denotes $u\bar{u}$, $d\bar{d}$, $s\bar{s}$, $c\bar{c}$, or $gg$.
 Process $\sigma(340.076)$ (fb)    $\sigma(341.598)$ (fb)    $\sigma(342.771)$ (fb)   $\varepsilon_{\text{tag}}(\%)$   $\varepsilon_{\text{hadron}}(\%)$
 $b\bar{b}$ 866.99 880.50 896.87 46.42 52.70
-$jj$ 4977.58 4945.68 4927.87 0.04 52.48
-$ZZ\to b\bar{b}b\bar{b}$ 16.17 16.01 15.86 0.31 0.31
-$ZH\to b\bar{b}b\bar{b}$ 12.14 11.88 11.70 0.31 0.31
-$ZZ\to b\bar{b}jj$ 116.42 115.11 114.19 1.64 2.21
-$ZH\to b\bar{b}jj$ 46.06 45.11 44.42 1.64 2.21
-$ZZ\to jjjj$ 209.21 207.11 205.38 0.00 5.19
-$ZH\to jjjj$ 8.60 8.42 8.29 0.00 5.19
-$WW\to jjjj$ 5263.55 5202.29 5157.25 0.00 5.19
-The cross section $\sigma(e^{+}e^{-}\to b\bar{b})$ is about 900 fb at around 340 GeV as shown in Fig.
-4, and the $\sum_{q=u,d,s,c,b}\sigma(e^{+}e^{-}\to q\bar{q})$ is about $6000$ fb at around 340 GeV.
-After our event selections, around 52% of the signal events passed the event selections, while nearly ${95}\%$ of the $WW/ZZ/ZH$ background events are rejected.
-Particularly, more than ${99}\%$ of the four-quark backgrounds are rejected in double $b$-tagged category.
-Once the efficiencies of all the involved processes have been established, the $R_{b}$ values at the three energy points can be obtained together with information of integrated luminosities, the cross sections, and the estimated numbers of signal and background events.
-A.4 Systematic uncertainties
-The systematic uncertainties, affecting the accuracy of $R_{b}$ measurements, arise from accelerator effects (e.g., $\sqrt{s}$ energy scale and energy spread), detector efficiencies (e.g., MC statistics, $\sqrt{s^{\prime}}$ and $p_{t}$ requirement), SM parameters ($\theta_{\rm W}$, $\alpha_{s}$, etc.), theoretical considerations (e.g., higher order corrections and ISR), and others (e.g., luminosity measurements and backgrounds from two-photon and $WWZ$ processes).
-All the systematic uncertainties are summarized in Table 4.
-Taking the $R_{b}$ measurement at the second energy point as an example, systematic uncertainties are discussed below in detail.
-The uncertainties related to accelerator include the $\sqrt{s}$ energy scale and energy spread.
-The uncertainty in the energy scale of $\sqrt{s_{2}}$ is estimated to be 2 MeV, the same as that at FCC-ee at 365 GeV [45].
-The uncertainty of the energy scale contributes $14.6\times 10^{-6}$ to $\Delta R_{b}$.
-The energy spread is 458.4 MeV for CEPC at the second energy point.
-A 10% uncertainty [15, 45] in the energy spread for CEPC or FCC-ee contributes an additional $2.7\times 10^{-6}$ to $\Delta R_{b}$.
-The uncertainties associated with the accelerator are different at the three energy points, consistent with the $R_{b}$ distribution in Fig.
-1.
-Notably, the first and third energy points corresponding to the minimum and maximum values of $R_{b}$ are sensitive to the energy spread, while the second energy point with the largest differential of $R_{b}$ over $m_{J_{t}}$ (very close to $\sqrt{s_{2}}$) is sensitive to the energy scale.
-The uncertainty associated with the detection efficiency in $R_{b}$ is ${\varepsilon_{\text{hadron}}^{udsc}}/{\varepsilon_{\text{tag}}^{b}}$, including MC statistics and the energy-momentum requirement, and is consistent with that in the $R_{b}$ measurement at the Z pole.
-In addition, the sizes of simulated events can be increased to achieve a statistical error less than $10^{-4}$, thus contributing $15.4\times 10^{-6}$ to $\Delta R_{b}$.
-We require $p_{t}>15~{}$ GeV and $\sqrt{s^{\prime}}>\sqrt{s}-10~{}$ GeV to select signal candidates, where the resolutions of $p_{t}$ and $\sqrt{s^{\prime}}$ also affect the detection efficiency.
-The jet energy and angular scale of the CEPC detector can be controlled within 0.5% and 0.01% [127].
-Thus, the precisions of $p_{t}$ and $\sqrt{s^{\prime}}$ can be controlled to within 0.10 GeV and 0.34 GeV, respectively, resulting in $\Delta R_{b}$ values of $9.1\times 10^{-6}$ and $6.7\times 10^{-6}$.
-The dominant backgrounds originate from four-fermion final states, such as $WW$, $ZZ$, and $ZH$ related processes.
-These processes are well understood theoretically and experimentally, and about 95% of such events are rejected by selection criteria.
-After collecting large data at the $Z/WW/ZH$ peaks at CEPC and FCC-ee, and taking into account the NNLO calculation, the relative uncertainty in these backgrounds will be less than $(\alpha_{s}/\pi)^{2}\sim 0.001$, thereby contributing $13.9\times 10^{-6}$ to $\Delta R_{b}$.
-The uncertainty in $\sin^{2}\theta_{\rm W}$, which can be achieved at $1\times 10^{-5}$ at CEPC [12, 45], contributes $6.0\times 10^{-6}$ to ${\Delta R_{b}}$.
-The uncertainty of $m_{b}$ in Eq.(15) contributes $0.2\times 10^{-6}$ to ${\Delta R_{b}}$.
-The uncertainty from $\alpha_{s}$ can be included in the higher order corrections.
-As shown in Eq.(A.2.2), the NNLO corrections will increase $R_{b}$ by $1.2\times 10^{-6}$.
-The uncertainty associated with ISR is canceled in $R_{b}$ measurement [43].
-The other sources include luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
-Their contributions to ${\Delta R_{b}}$ are small, with a combined estimation of $<3.0\times 10^{-6}$.
-Assuming all sources of systematic uncertainties in $R_{b}$ are independent, their contributions are combined in quadrature to result in a total systematic uncertainty of $R_{b}$ at the second energy point:
-$({\Delta R_{b}})_{\text{syst.}}=28.6\times 10^{-6}.$ (26)
-As shown in Fig.
-1, the difference in $R_{b}$ between the first and second points is $(5567.6\pm 1002.1_{\rm stat.})\times 10^{-6}$, compared to SM theoretical prediction of $-41.9\times 10^{-6}$.
-The total systematic uncertainty in the energy scale and energy spread is ±$\pm 39.0\times 10^{-6}$, which means that the difference in $R_{b}$ is $5.6\sigma$.
-Other systematic uncertainties in the detection efficiency, backgrounds, etc.
-do not change this difference.
-Table 4: Uncertainties of $R_{b}^{\rm PEX}$ are given in a unit of $10^{-6}$.
-The “Other” category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
-As a comparison, the pseudo-experiment $R_{b}^{\rm PEX}$ and SM prediction $R_{b}^{\rm SM}$ (in a unit of $10^{-6}$) are also provided here.
-[…]
+$jj$ 4,977.58 4,945.68 4,927.87 0.04 52.48
+$ZZ/ZH\to b\bar{b}b\bar{b}$ $16.17/12.14$ $16.01/11.88$ $15.86/11.70$ 0.31 0.31
+$ZZ/ZH\to b\bar{b}jj$ $116.42/46.06$ $115.11/45.11$ $114.19/44.42$ 1.64 2.21
+$ZZ/ZH/WW\to jjjj$   $209.21/8.60/5,\!263.55$   $207.11/8.42/5,\!202.29$   $205.38/8.29/5,\!157.25$ 0.00 5.19
+Table 4: Uncertainties of $R_{b}^{\rm PEX}$ are given in a unit of $10^{-6}$. The “Other" category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
+Source 340.076 GeV    341.598 GeV    342.771 GeV
 Energy scale 0.5 14.6 0.1
@@
 $\sqrt{s^{\prime}}$ requirement 6.4 6.7 7.0
-Background 13.5 13.9 14.2
+Backgrounds 13.5 13.9 14.2
 SM parameters 5.0 6.0 6.8
 Higher order corrections 1.2 1.2 1.2
-Other $<3.0$ $<3.0$ $<3.0$
+Other $3.0$ $3.0$ $3.0$
 ${\rm Sum~{}of~{}syst.}$ 43.0 28.6 64.0
-${\rm Stat.}$ 884.4 471.2 1859.3
-${\rm Total}$ 885.5 472.1 1860.4
-$R_{b}^{{\rm PEX}}(\sqrt{s_{i}})$     146970.0     152537.6     156589.0
-$R_{b}^{{\rm SM}}(\sqrt{s_{i}})$ 151539.9 151498.0 151466.1
+${\rm Stat.}$ 884.4 471.2 1,859.3
+${\rm Total}$ 885.5 472.1 1,860.4